### v2

# DPLib: A Standard Benchmark Library for Distributed Power System Analysis and Optimization

Abstract
DPLib is an open-source MATLAB-based benchmark library created to support research and development in distributed and decentralized power system analysis and optimization. Distributed and decentralized methods offer scalability, privacy preservation, and resilience to single points of failure, making them increasingly important for modern power systems. However, unlike centralized tools such as MATPOWER, no general-purpose, reproducible data library package currently exists for distributed power system studies. DPLib, available at GitHub, fills this gap by providing a standard power system library featuring over 20 multi-region benchmark test cases of varying sizes, along with a graph-based partitioning toolkit that decomposes any MATPOWER test system into multiple electrically coherent regions. The partitioning toolkit, an easy-to-use MATLAB code, generates standardized .mat and .m files, along with region visualizations for intuitive understanding. We also provide modular, easy-to-use distributed optimal power flow (OPF) solvers: an alternating direction method of multipliers(ADMM)-based DC-OPF solver implemented in YALMIP, and an ADMM-based AC-OPF solver leveraging IPOPT. These solvers validate the generated test systems for distributed optimization applications. Numerical results validate the generated test cases, establishing DPLib as a foundation for reproducible distributed power system research.
Index Terms:
Distributed power grid test systems, standard benchmark library, MATPOWER, distributed optimization.
I Introduction
Centralized frameworks have long served as the dominant approach for modeling and solving large-scale power system problems, including optimal power flow (OPF) and security-constrained unit commitment. A key enabler behind the success and widespread adoption of centralized optimization and analysis in power systems is the availability of standardized benchmark test cases and libraries. Among these, MATPOWER has emerged as a widely accepted benchmark tool, offering consistent, well-structured data for centralized OPF studies  [1]. Its role in enabling reproducibility, fair performance comparisons, and methodological development has been fundamental to the evolution of centralized power system optimization and analysis.
For centralized OPF studies, for instance, [2] proposed convex relaxations for AC OPF and validated their results on MATPOWER test cases. [3] provided a comprehensive review of relaxation techniques for OPF, with all benchmarks sourced from MATPOWER. [4] introduced QC relaxation for OPF using MATPOWER data. More recently, [5] developed robust optimization methods for corrective control under uncertainty, relying on MATPOWER cases for validation. [6] proposed the NESTA archive to enrich MATPOWER with more realistic data, yet retained full MATPOWER compatibility. These works and many more underscore the importance of a common, reliable data standard—facilitated by MATPOWER—that allows researchers to validate and compare centralized optimization algorithms under consistent and reproducible conditions.
In the last decade, distributed and decentralized computation has gained significant momentum in the power systems community, especially in applications such as distributed OPF, voltage control, state estimation, and demand-side management [7, 8, 9]. These include studies that apply alternating direction method of multipliers (ADMM) decomposition [8, 10], distributed voltage control [11], parallel and distributed OPF [12, 13], consensus-based optimization [14], and cooperative control frameworks [15, 16]. Additional contributions include scalable distributed frameworks for transmission networks [17], graph-based coordination strategies [7], and distributed control in microgrids [18]. These methods offer inherent advantages in scalability, privacy preservation, and resilience to single points of failure.
However, as the literature on distributed techniques expands, a critical gap remains: the lack of unified, publicly available benchmark test cases tailored for distributed power system studies. Without such a standard library, researchers must often construct ad hoc datasets or apply arbitrary partitioning schemes, which hampers reproducibility and limits the comparability of different distributed algorithms. There is a strong need for a publicly available and community-accepted library of distributed test cases that allows researchers to validate, compare, and benchmark their distributed optimization or control methods under a common, well-defined framework.
In response to this need, we develop a distributed power system library, DPLib, that addresses the lack of standardized resources for distributed power system studies by providing a comprehensive library of over 20 benchmark test cases, ranging from 73 to 9241 buses. These multi-region systems are generated using the DPLib graph-based partitioning toolkit that ensures electrical coherence and captures the corresponding multi-regional network topology. Each case includes the necessary structural and connectivity data to support distributed analysis. All test cases have been verified using distributed OPF problems to demonstrate their practicality and confirm their suitability for reproducible and scalable distributed optimization.
In addition to a standard library, a key requirement is a general and user-friendly partitioning tool capable of processing any standard MATPOWER case and dividing it into electrically coherent subregions based on user-defined parameters, such as the number of regions. An effective tool should preserve the integrity of the original network, identify and document inter-regional tie-lines, and produce output in a structured format consistent with centralized MATPOWER cases. This functionality enables users to generate customized test cases while maintaining compatibility with a standard data structure. The DPLib graph-based partitioning toolkit fulfills all of these requirements.
While the distributed library includes several pre-partitioned and validated datasets, researchers may also wish to generate and test their customized partitions using the DPLib graph-based toolkit. To support this, DPLib provides ADMM-based solvers for both DC OPF and AC OPF that are fully compatible with the partitioned data structure produced by the tool. These solvers take the partitioned datasets as input, iteratively solve the regional subproblems, enforce tie-line consistency through Lagrangian multipliers, and monitor convergence metrics, enabling reliable validation of user-defined partitions.
The main contribution of this work is the introduction of DPLib, a unified framework that establishes the first standard distributed data library for power system applications. Besides the library, DPLib consists of three core tools: a graph-based partitioning tool that decomposes any MATPOWER case into multiple regions using spectral clustering and k-means, enabling the creation of distributed datasets; a modular ADMM-based distributed DC OPF solver implemented in YALMIP; and a distributed AC OPF solver using ADMM and IPOPT. The DC OPF and AC OPF solvers are designed to validate and benchmark the partitioned datasets under realistic operating conditions. Together, these tools support reproducibility, scalability, and algorithm testing for distributed power system applications, addressing a key gap in current research infrastructure.
The remainder of the paper is organized as follows. Section II introduces the graph-based modeling of power networks and details the DPLib partitioning toolkit. Section III describes the ADMM-based DC and AC OPF solvers used to validate partitioned datasets. Section IV presents the benchmark case studies generated and tested using DPLib. Finally, Section V concludes the paper.
II Graph-Based Partitioning and Toolkit Description
In many power systems studies, realistic standard test cases in the MATPOWER format are commonly used for testing and validation. This practice ensures that developed methods are evaluated on realistic cases that reflect the complexity of actual power systems. To maintain this standard, researchers working on distributed algorithms should also rely on realistic standard test systems and datasets rather than arbitrarily constructed or manually partitioned grids. In practice, real power networks often exhibit a naturally regional structure. For example, a network covering several regions will typically have strong internal connectivity within each region and only a few tie-lines connecting areas to one another. This results in regions that are strongly connected internally but weakly connected to the other areas. A proper partitioning method should be able to observe these loosely connected boundaries. We use graph-based partitioning to provide a principled and effective framework for capturing such structures and dividing the network in a way that aligns with real systems’ physical and topological characteristics.
II-A Power Network Modeling and Spectral Clustering
To enable distributed optimization and control, a centralized network must first be divided into smaller regions that remain physically meaningful and electrically coherent. This partitioning process starts by modeling the power system as an undirected graph. We represent the power network as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where:
  • $\mathcal{V}$ is the set of buses (nodes), as defined in the original MATPOWER case file.
  • $\mathcal{E}$ is the set of transmission lines (edges), each connecting two buses.
The graph is undirected because the power flow direction does not impact the structure for partitioning purposes—only a connection between buses matters.
A path in a graph is a sequence of nodes such that an edge connects each consecutive pair of nodes in the sequence. A connected graph is an undirected graph with a path between every pair of nodes. This means it is possible to travel from any node to any other by following a sequence of edges in the graph. To describe the structure of $\mathcal{G}$, we construct the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $n=|\mathcal{V}|$ is the total number of buses. The entries of this matrix are defined as:
$A_{ij}=\begin{cases}1,&\text{if there is a line connecting bus }i\text{ and bus }j,\\ 0,&\text{otherwise}.\end{cases}$
Because the graph is undirected, $A$ is symmetric. Next, we define the degree matrix $D\in\mathbb{R}^{n\times n}$, a diagonal matrix where each diagonal entry $D_{ii}$ equals the number of lines connected to bus $i$:
$D_{ii}=\sum_{j=1}^{n}A_{ij}.$
Using these two matrices, we define the graph Laplacian $L=D-A.$ The Laplacian matrix $L$ encodes the power network connectivity and topology. For a connected graph, $L$ is symmetric and positive semi-definite, meaning all its eigenvalues are non-negative. Importantly, the smallest eigenvalue of $L$ is always zero, corresponding to the trivial uniform eigenvector.
To divide the system into $k$ electrically coherent regions, we use spectral clustering, which leverages the properties of $L$. The first step is to compute the eigenvalues and eigenvectors of $L$. Among these, the $k$ smallest nonzero eigenvalues are selected. Each eigenvalue is associated with an eigenvector that varies smoothly across the graph, meaning that buses that are strongly connected in the physical system will have similar eigenvalues. Therefore, these eigenvectors reflect the underlying structure of the network.
We form a matrix $U\in\mathbb{R}^{n\times k}$ by placing the $k$ selected eigenvectors as columns. Each row $u_{i}\in\mathbb{R}^{k}$ corresponds to a bus and represents it in a new $k$-dimensional spectral space, where geometrical proximity reflects topological connectivity.
To assign buses into one of $k$ regions, we apply the k-means clustering algorithm to the rows of $U$. This unsupervised learning algorithm operates by:
  • Randomly initializing $k$ centroids (in our case, from the rows of $U$);
  • Assigning each bus to the nearest centroid based on Euclidean distance;
  • Recalculating the centroid of each cluster;
  • Repeating the assignment and centroid update steps until convergence.
The clustering result yields a partition of all buses into $k$ regions. Each cluster groups strongly connected buses, leading to electrically coherent and balanced regions. After forming the clusters, inter-regional tie-lines (lines that connect buses belonging to different clusters) are identified. To improve quality, the algorithm may repeat the clustering multiple times and select the configuration that minimizes the number of inter-regional tie-lines.
This spectral clustering approach provides a systematic and scalable method for decomposing any MATPOWER case into meaningful regions. It aligns well with practical expectations since most real-world partitioned power networks are formed based on geographic or electrical proximity.
A tool that can automatically convert MATPOWER test cases into $k$ well-structured regions is essential for distributed power system studies. The DPLib partitioning toolkit fills this gap by generating standardized multi-region datasets from any MATPOWER case, supporting reproducibility and scalability.
II-B DPLib Partitioning Toolkit
The DPLib graph-based partitioning toolkit implements the method described above and automates the generation of partitioned MATPOWER cases. The partitioning toolkit is publicly available at GitHub, including the MATLAB implementation file named partitioning_code.m. The code uses MATLAB R2020a or newer, along with the MATPOWER toolbox. Upon running the script, the user is prompted to enter the name of a valid MATPOWER case file (for example, case118) and specify the desired number of regions into which the system should be divided.
Internally, the code loads the MATPOWER case and checks whether the bus numbers are sequential. If not, it renumbers the buses from one to $n$ and updates all corresponding references in the branch and generator matrices. It then constructs the undirected adjacency matrix based on the network’s branch connectivity and forms the corresponding graph Laplacian. The eigenvectors of the Laplacian are computed and used as features for $k$-means clustering. The clustering process is repeated multiple times to ensure optimality, and the configuration with the lowest number of inter-regional tie-lines is selected.
For each candidate partitioning, regional subcases are formed by extracting internal buses, branches, generators, and generation cost data for each region. The script ensures that each region has at least one generator. Buses are then renumbered locally within each region to maintain consistency. Inter-regional tie-lines are identified as any branch connecting buses in different regions and stored.
Once a feasible partitioning is found, the code saves the result in multiple formats. A MATLAB .mat file is created, containing all regional subcases under separate variables (e.g., mpc_regionR1, mpc_regionR2, etc.), along with a comprehensive table of all inter-regional tie-lines labeled by their originating and receiving regions and complete branch data. A corresponding .m file is also generated, exporting the same structures in a human-readable script format suitable for inspection or direct use in other codes. Additionally, the code creates a graphical visualization of the partitioned system, where each region is displayed as a labeled node in a circular layout, and tie-lines are represented as edges connecting the areas. The number of tie-lines between each pair of regions is also annotated.
During execution, the code prints detailed messages to the MATLAB command window. These include whether bus renumbering was necessary, confirmation of full graph connectivity, the number of generated regions and tie-lines, the region assignment of the reference (slack) bus, and whether all buses, branches, and generators are fully included in the partitioned result. It also confirms the successful saving of the generated .mat, .m, and .png files. The algorithm of the partitioning toolkit, illustrating the complete data processing pipeline, is provided in 1.
Algorithm 1 DPLib partitioning toolkit
1:User Input:
  • Name of the original MATPOWER case (e.g., case118)
  • Number of desired regions $k$
2: Load the MATPOWER case file
3: Check bus numbering; if non-sequential, renumber from 1 to $n$ and update all references
4: Construct adjacency matrix $A$ from the branch data.
5: Construct degree matrix $D$ with $D_{ii}=\sum_{j}A_{ij}$
6: Form graph Laplacian $L=D-A$
7: Compute the first $k$ nonzero eigenvectors of $L$
8: Stack eigenvectors into matrix $U\in\mathbb{R}^{n\times k}$
9: Apply $k$-means clustering to the rows of $U$
10: Repeat clustering for multiple iterations to minimize the number of tie-lines
11:for each region (cluster) do
12:  Extract internal buses, branches, generators, and cost data
13:  Ensure that at least one generator exists in each region
14:  Locally renumber buses
15:  Identify inter-regional tie-lines
16:end for
17: Save output in the following formats:
  • .mat file with region-wise MATPOWER data
  • .m script file exporting the same data
  • .png visualization showing regions and tie-lines
III Validating DPLib Test Systems Using OPF
To ensure that the generated multi-region test cases are electrically coherent and suitable for distributed power system studies, it is essential to validate them through power system analysis. Although the DPLib partitioning toolkit generates distributed variants of MATPOWER test systems applicable to a wide range of power system studies, including control, state estimation, or stability analysis, this section evaluates these test systems for distributed OPF applications. OPF has been selected as a representative benchmarking problem because of its maturity, availability of centralized references, and compatibility with distributed solvers due to over a decade of research in this area. We have developed general ADMM-based solvers for DC and AC OPF problems that operate directly on the test cases produced by the DPLib partitioning toolkit. These solvers are designed to validate the feasibility, consistency, and convergence of partitioned datasets, demonstrating the suitability of these test systems. Nevertheless, users may use alternative solvers or utilize these provided tools to verify and explore their own partitioned datasets.
III-A Distributed OPF Formulation
We formulate the OPF problem in both DC and AC settings for each region $r$. Each region solves a local OPF problem while maintaining consistency at the boundaries via duplicated (virtual) variables and tie-line constraints. These virtual variables (buses) enable each region to model tie-line interactions locally without requiring direct enforcement of global constraints. The overall structure is illustrated in Fig. 1(a), where a two-region power grid is shown with inter-regional tie-lines in red. Fig. 1(b) further illustrates how virtual buses (shown in red) are added to represent shared variables.
(a) Grid with two regions and red lines are inter-regional transmission tie-lines
(b) Decomposition into two regions with virtual buses shown in red
Figure 1: Illustration of a partitioned power system
III-A1 DC OPF Formulation
The DC OPF for region $r$ is:
$\min_{\boldsymbol{x}^{r}}\sum_{i\in\Omega^{r}_{g}}f(p^{g}_{i})$ (1a)
s.t.
$\qquad\quad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (1b)
$\qquad\quad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\frac{\theta_{i}-\theta_{j}}{x_{ij}}+\sum_{j\in\Omega^{r,\textrm{in}}_{i}}\frac{\theta_{i}-\tilde{\theta}_{j}}{x_{ij}},\quad\forall i\in\Omega^{r}_{b}$ (1c)
$\qquad\quad\underline{p}_{ij}\leqslant p_{ij}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (1d)
$\qquad\quad\theta_{\text{ref }}=0,$ (1e)
$\qquad\quad\theta_{i}=\tilde{\theta}_{i},\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (1f)
$\boldsymbol{x}^{r}=[\boldsymbol{p}^{g\top},\boldsymbol{\theta}^{\top}]\in\mathbb{R}^{n^{r}_{g}+n^{r}_{b}}$ is the decision vector for region $r$, including generation $p^{g}_{i}$ and voltage angles $\theta_{i}$. Equation (1a) minimizes the local cost $f(p^{g}_{i})$. Constraints (1b)–(1c) enforce generator limits and power balance using demand $p^{d}_{i}$, line reactance $x_{ij}$, and virtual angles $\tilde{\theta}_{j}$. Constraint (1d) bounds internal and tie-line flows $p_{ij}$, and (1f) ensures consistency between real and virtual bus angles.
III-A2 AC OPF Formulation
The AC OPF for region $r$ is:
$\min_{\boldsymbol{x}^{r}}\sum_{i\in\Omega^{r}_{g}}f(p^{g}_{i})$ (2a)
s.t.
$\qquad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\underline{q}_{i}^{g}\leqslant q^{g}_{i}\leqslant\bar{q}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (2b)
$\qquad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(p_{ij,f}+p_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}\tilde{p}_{ij},\quad\forall i\in\Omega^{r}_{b}$ (2c)
$\qquad q^{g}_{i}-q^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(q_{ij,f}+q_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}\tilde{q}_{ij},\quad\forall i\in\Omega^{r}_{b}$ (2d)
$\qquad p_{ij,f}=G_{ij,1}v_{i}^{2}+G_{ij,2}v_{i}v_{j}\cos(\theta_{i}-\theta_{j})$
$\qquad\qquad\quad+B_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (2e)
$\qquad q_{ij,f}=-B_{ij,1}v_{i}^{2}-B_{ij,2}v_{i}v_{j}\cos(\theta_{i}-\theta_{j})$
$\qquad\qquad\quad+G_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (2f)
$\qquad p_{ij,t}=G_{ij,4}v_{j}^{2}+G_{ij,3}v_{i}v_{j}\cos(\theta_{j}-\theta_{i})$
$\qquad\qquad\quad+B_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (2g)
$\qquad q_{ij,t}=-B_{ij,4}v_{j}^{2}-B_{ij,3}v_{i}v_{j}\cos(\theta_{j}-\theta_{i})$
$\qquad\qquad\quad+G_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (2h)
$\qquad p_{ij,f}^{2}+q_{ij,f}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2i)
$\qquad p_{ij,t}^{2}+q_{ij,t}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2j)
$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\theta_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2k)
$\qquad\theta_{\text{ref}}=0$ (2l)
$\qquad\theta_{i}=\tilde{\theta}_{i},\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (2m)
$\qquad v_{i}=\tilde{v}_{i},\quad\tilde{v}_{j}=v_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (2n)
The decision vector is $\boldsymbol{x}^{r}=[\boldsymbol{p}^{g\top},\boldsymbol{q}^{g\top},\boldsymbol{\theta}^{\top},\boldsymbol{v}^{\top}]$, including generation $(p^{g}_{i},q^{g}_{i})$, angles $\theta_{i}$, and voltages $v_{i}$ at region $r$. Constraints (2b)–(2d) enforce generator and nodal power limits; (2e)–(2h) define AC flows using admittance terms $G_{ij,k},B_{ij,k}$. Limits on flows and angle differences are imposed in (2i)–(2k). Reference angle (2l) and virtual bus equalities (2m)–(2n) ensure boundary consistency.
III-A3 ADMM Coordination
To ensure consistency between neighboring regions, we use ADMM to coordinate virtual variables [10]. For example, the voltage angle consistency is enforced via:
$L_{\rho}(\theta,\tilde{\theta},\lambda)=f(p^{g})+\lambda^{\top}(\theta-\tilde{\theta})+\frac{\rho}{2}\|\theta-\tilde{\theta}\|^{2},$ (3)
where $\lambda$ are the dual variables and $\rho$ is the penalty parameter. Similar terms are added for voltage magnitudes in AC OPF.
Each ADMM iteration proceeds as:
  1. 1.
    Solve each region’s local OPF using the current dual variables.
  2. 2.
    Update dual variables based on tie-line consistency.
  3. 3.
    Check primal residuals and stop if below a threshold.
In general, coordination across regions is achieved using the ADMM. The distributed DC OPF and AC OPF formulations in (1) and (2) incorporate local objectives and tie-line consistency via augmented Lagrangian. Subproblems are solved locally at each iteration, dual variables are updated, and convergence is checked via the worst primal residual.
III-B General ADMM-Based OPF Solvers
We have developed two open-source general-purpose solvers for distributed OPF problems, one for DC OPF and another for AC OPF. Both solvers adopt ADMM to coordinate regional solutions and ensure consistency across tie-lines.
III-B1 DPLib DC OPF Solver
The DPLib DC OPF solves the OPF problem using ADMM by optimizing each region independently and enforcing agreement on shared variables (tie-line voltage angles) through iterative dual updates. The solver is available at GitHub, including the MATLAB implementation file named DistDCOPF.m. Designed for partitioned MATPOWER cases, it requires MATLAB R2020a or newer with MATPOWER and YALMIP installed. Upon execution, the user provides the partitioned .mat file, ADMM penalty parameter $\rho$, convergence threshold, maximum number of iterations, and optionally the centralized OPF cost for benchmarking via optimality gap.
After receiving user inputs, the code loads all regional and tie-line data, augments each region’s branch matrix with inter-regional tie-lines, and initializes bus voltage angle variables and indexing. The core ADMM loop then solves each region’s DC OPF using YALMIP, minimizing generation cost under local constraints (e.g., power balance, generator limits) while adding consistency penalties on shared tie-line angles as modeled in (1). A helper function handles the regional problem formulation and solution.
After each iteration, the code evaluates tie-line angle consistency and updates dual variables accordingly. Primal residuals are recorded, and convergence is declared if the maximum error drops below the user-defined threshold; otherwise, the process continues until the iteration limit. The optimality gap is computed and plotted alongside residuals if a centralized cost is provided. Progress and convergence status are printed in the command window.
The DPLib DC OPF solver supports any number of regions and is tailored for use with partitioned .mat files, enabling reproducible distributed OPF testing. Algorithm 2 summarizes its workflow.
Algorithm 2 DPLib OPF solver
1:Input: Partitioned MATPOWER data, ADMM parameters $\rho$, convergence tolerance $\epsilon$, maximum iterations
2: Load the partitioned case data for all regions
3: Augment each region with its corresponding tie-line data
4: Initialize region variables including local decision variables and dual variables
5: Retrieve ADMM parameters and consistency constraints
6: Identify and prepare interface buses at region boundaries
7:for each ADMM iteration do
8:  Solve local OPF for each region independently
9:  Update dual variables based on consistency violations at bordering buses
10:  if convergence criteria are met then
11:   break
12:  end if
13:end for
14: Visualize and save final solution
III-B2 DPLib AC OPF Solver
The DPLib AC OPF solver implements a distributed ADMM-based solver in MATLAB for any power system preprocessed using the DPLib partitioning tool. It enables large-scale distributed analysis, where each region independently solves its local AC OPF while coordinating inter-regional tie-lines. The solver developed in this work is publicly available at GitHub, including the MATLAB implementation file named DistACOPF.m. To run the code, the user must have MATLAB R2020a or later, with MATPOWER and IPOPT properly configured. Upon launch, the code prompts for the partitioned .mat file, ADMM penalty parameter $\rho$, convergence threshold, maximum number of iterations, and optionally the centralized OPF cost for benchmarking via optimality gap.
After loading the partitioned data, the code augments each region’s branch matrix with its tie-lines and initializes region-specific variables (e.g., bus counts, initial voltage values). The main ADMM loop iteratively solves each region’s local AC OPF via the helper function Acopf, based on (2), updating voltage magnitudes and angles. Dual variables are adjusted using consistency gradients to enforce agreement across tie-lines. Convergence is declared if the worst primal residual falls below the threshold. The optimality gap is computed per iteration if a centralized cost is provided. Final plots show convergence trends, and status messages are printed throughout. The overall process flow of the DPLib AC OPF solver is presented in algorithm 2.
IV DPLib: Benchmark Test Cases
Distributed versions of many widely used MATPOWER benchmark systems are generated. Each test case is processed using the DPLib graph-based partitioning tool and tested using the DPLib ADMM-based DC and AC OPF solvers described earlier. Across all systems, both solvers successfully converge to feasible solutions. All partitioned cases and solver outputs are publicly available as part of the DPLib at GitHub. In the following subsections, we explain some of the distributed test cases generated and tested as benchmarks in DPLib.
In addition to evaluating convergence using the worst-case primal residual and the optimality gap, further validation tests are performed on the generated multi-region systems to verify their consistency with physical laws and operational constraints of the power grid. Specifically, we verified that upon the convergence of the ADMM iterations, the total power generation across all regions closely matches the total system demand, accounting for minor discrepancies due to modeling approximations and convergence thresholds. Furthermore, we confirmed that active power flows across each inter-regional tie-line are consistent from both sides of the interface. All tested datasets successfully passed these validation checks, further demonstrating the correctness and reliability of the generated partitioned cases.
To facilitate evaluation of the benchmark test cases using OPF solvers, all primal variables were initialized with cold-start conditions: voltage angles were set to zero and voltage magnitudes to one per unit. The dual variables, corresponding to Lagrange multipliers associated with consistency constraints, were initialized to one. Penalty parameters $\rho$ in the ADMM-based solvers were selected case by case to ensure convergence.
IV-A Case: pglib_opf_case200_tamu
The pglib_opf_case200_activ system is a synthetic 200-bus benchmark from the IEEE PES Power Grid Library for optimal power flow studies. Geographically situated in the central region of Illinois, this case models a 230/115 kV transmission network designed to serve a load distribution that mimics the population profile of its geographic footprint. The transmission system was generated following the methods outlined in [19], ensuring statistical similarity to real-world grids without modeling any actual transmission lines. Developed as part of the U.S. ARPA-E GRID DATA research project, the system contains no critical energy infrastructure information and was created using PowerWorld Simulator, then converted to MATPOWER format using version 6. The dataset is publicly available through PGLib and licensed under the Creative Commons Attribution 4.0 International License.
The pglib_opf_case200_tamu test case was partitioned using our graph-based spectral clustering method, which ensured full bus inclusion and topological connectivity. The best clustering configuration identified by the algorithm consisted of three electrically coherent regions, interconnected through 10 tie-lines, within 10 iterations. During the partitioning, the reference bus, bus 189, was assigned to Region 2 (mpc_regionR2). No buses were omitted, and the total number of buses across all regions matches the original MATPOWER dataset, with 200 buses, 245 branches, and 49 generators. The consistency of branch count, generator count, and overall system topology confirms the fidelity of the partitioning process. The final multi-region topology has been visualized and is shown in Fig. 2. The corresponding partitioned files are saved as pglib_opf_case200_tamu_3regions.mat and pglib_opf_case200_tamu_3regions.m. Both distributed DC OPF and AC OPF solvers have been executed on this partitioned data, and their convergence trends are illustrated through the worst primal residuals presented in Fig. 3 with $\rho=1000$.
Figure 2: Topology of pglib_opf_case200_tamu partitioned into three regions with 10 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 3: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case200_tamu.
IV-B Case: pglib_opf_case300_ieee
The pglib_opf_case300_ieee system is a widely used benchmark from the IEEE PES Power Grid Library for optimal power flow studies. This test case represents the IEEE 300-bus system, which was initially provided in the IEEE Common Data Format and converted into MATPOWER format on September 20, 2004, using the cdf2matp utility. The dataset was sourced from the University of Washington’s power systems test case archive. The system corresponds to typical operational conditions, with a base power of 100 MVA and includes detailed modeling of generation, loads, and transmission elements.
Using our graph-based spectral clustering approach, we have partitioned the pglib_opf_case300_ieee system into three electrically coherent regions. Before clustering, the original bus numbering in the MATPOWER file was reordered to ensure consistent indexing. The graph topology of the network was verified to be fully connected, allowing for a valid partitioning process. The algorithm identified the best clustering configuration in 10 iterations, resulting in 3 regions interconnected through 12 tie-lines. The system’s slack bus, originally defined at bus 257, was successfully assigned to Region 1 (denoted as mpc_regionR1). After partitioning, the integrity of the dataset was confirmed: all 300 buses, 411 branches, and 69 generators were fully included in the regional breakdown, with no data omitted. The final partitioned system has been saved as both a .mat file and a script-based .m file named pglib_opf_case300_ieee_3regions. A visual representation of the multi-regional topology is shown in Fig. 4. The distributed DC OPF and AC OPF solvers were successfully executed on this dataset, and the convergence behavior is demonstrated through the worst primal residual plots shown in Figs. 5(a) and 5(b) with $\rho=1000$.
Figure 4: Topology of pglib_opf_case300_ieee partitioned into three regions with 12 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 5: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case300_ieee.
IV-C Case: pglib_opf_case500_tamu
The pglib_opf_case500_tamu case is a synthetic 500-bus power system in the IEEE PES Power Grid Library for OPF studies. This case was developed as part of the ARPA-E GRID DATA research initiative and is geographically based in the northwestern region of South Carolina. It incorporates a transmission network operating at 345 kV and 138 kV voltage levels, constructed to supply power to a synthetic load distribution that mirrors the population density of the modeled region. The system does not represent any real grid infrastructure but is created using statistical modeling techniques described in [19], ensuring structural similarities to real transmission systems while maintaining confidentiality.
Using our graph-based spectral clustering framework, the pglib_opf_case500_tamu benchmark system was partitioned into five electrically coherent regions. Before clustering, the bus numbers from the original MATPOWER file were analyzed to preserve consistent indexing. The graph representation of the system was confirmed to be fully connected, ensuring the validity of the partitioning procedure. The spectral clustering algorithm converged in 10 iterations and identified an optimal structure of 5 regions interconnected through 16 tie-lines. The reference bus, originally defined as bus 17, was successfully assigned to Region 2 (denoted as mpc_regionR2). Post-partitioning checks confirmed that all 500 buses, 597 branches, and 90 generators were preserved without omission or duplication. The partitioned data has been saved in both .mat and script-based .m formats under the name pglib_opf_case500_tamu_5regions. A visual illustration of the multi-regional layout is presented in Fig. 6. Distributed DC OPF and AC OPF solvers were successfully executed for this case, and the convergence profiles are illustrated by the worst primal residual curves in Figs. 7(a) and 7(b) with $\rho=1000$.
Figure 6: Topology of pglib_opf_case500_tamu partitioned into five regions with 16 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 7: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case500_tamu.
IV-D Case: pglib_opf_case1354_pegase
The pglib_opf_case1354_pegase dataset is a detailed test case that emulates the size and complexity of the European high-voltage transmission network. It includes 1,354 buses, 260 generators, and 1,991 branches operating at 220 and 380 kV voltage levels. While the data is fictitious and does not reflect real-world infrastructure, it is highly suitable for validating power system optimization tools and techniques. This benchmark was developed as part of the Pan European Grid Advanced Simulation and State Estimation (PEGASE) project. It incorporates unique modeling features such as asymmetric shunt elements and generator aggregations.
The pglib_opf_case1354_pegase benchmark system was partitioned into eight electrically coherent regions using our graph-based spectral clustering method. The original bus indices from the MATPOWER file were reordered to ensure consistency during partitioning. The graph representation of the power network was verified to be fully connected, ensuring that a valid and complete partitioning could be achieved. The clustering algorithm successfully identified an optimal partitioning configuration within 10 iterations, resulting in 8 regions interconnected through 62 tie-lines. The system’s reference bus, which is located at bus 640 in the original data, was assigned to Region 1 (denoted as mpc_regionR1). The integrity of the data was verified after partitioning: all 1354 buses, 1991 branches, and 260 generators were accurately distributed across the regions with no missing components. The partitioned data files were saved in both .mat and .m formats under the name pglib_opf_case1354_pegase_8regions. A graphical representation of the resulting multi-regional topology is provided in Fig. 8. Both distributed DC OPF and AC OPF solvers were applied to this case study, with convergence behavior illustrated via worst primal residual plots in Figs. 9(a) and 9(b) with $\rho=1000$.
Figure 8: Topology of pglib_opf_case1354_pegase partitioned into eight regions with 62 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 9: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case1354_pegase.
IV-E Case: pglib_opf_case2383wp_k
The pglib_opf_case2383wp_k dataset models the Polish national transmission grid under peak load conditions during the winter of 1999–2000. It includes a comprehensive representation of 400, 220, and 110 kV networks and originates from a reduced version of the broader 7500+ bus European UCTE system. To simplify the model while preserving essential characteristics, tie lines connecting to neighboring countries were replaced by artificial load or generator buses. Generator aggregation and cost approximations were applied to reflect dispatchability in the Polish energy market.
Using the proposed graph-based spectral clustering method, the pglib_opf_case2383wp_k benchmark system was partitioned into twelve electrically meaningful regions. The underlying graph structure of the power system was verified to be fully connected, confirming the feasibility of multi-regional partitioning. The algorithm determined the optimal clustering within 10 iterations, producing 12 regions linked by 99 inter-regional tie-lines. The system’s slack bus, originally defined at bus 18 in the MATPOWER case, was assigned to Region 4 (referred to as mpc_regionR4).
Following partitioning, a full consistency check verified that all 2383 buses, 2896 branches, and 327 generators were preserved and correctly distributed across the regional datasets. No components were omitted during the process. The results were stored as both a .mat file and an executable .m script, saved under the name pglib_opf_case2383wp_k_12regions. The partitioned network topology is visually represented in Fig. 10. The distributed DC OPF and AC OPF solvers are implemented and their convergence behavior is shown through the worst primal residual plots in Figs. 11(a) and 11(b) with $\rho=2000$.
Figure 10: Topology of pglib_opf_case2383wp_k partitioned into twelve regions with 99 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 11: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case2383wp_k.
IV-F Case: pglib_opf_case3012wp_k
The pglib_opf_case3012wp_k dataset provides a detailed model of the Polish transmission grid during the winter 2007–2008 evening peak demand period. It represents the operational state of the 400, 220, and 110 kV networks and is a high-fidelity snapshot of the Polish system during that time. As part of the Power Grid Library benchmark suite, it is widely used for testing and validating optimization algorithms in realistic, large-scale scenarios.
Using the proposed graph-based spectral clustering framework, the pglib_opf_case3012wp_k benchmark system was partitioned into twelve electrically coherent regions. Before partitioning, the original MATPOWER bus numbers were reordered to enable consistent indexing throughout the graph representation. The full graph connectivity was verified to ensure that all buses could be properly included in the clustering process. Within 10 iterations, the partitioning algorithm identified an optimal configuration that divides the system into 12 regions connected by 100 inter-regional tie-lines. The slack bus, located initially at bus 37 in the MATPOWER data, was assigned to Region 12 (denoted as mpc_regionR12).
Following the partitioning process, a verification step confirmed that the distributed representation included all 3012 buses, 3572 branches, and 502 generators. These components were preserved without omission or duplication. The resulting data was saved both as a .mat file and as a script-based .m file under the name pglib_opf_case3012wp_k_12regions. A visual overview of the final partitioned topology is shown in Fig. 12. Using this distributed dataset, the DC OPF and AC OPF solvers were executed successfully, with their convergence behavior illustrated by the worst primal residual plots shown in Figs. 13(a) and 13(b) with $\rho=2000$.
Figure 12: Topology of pglib_opf_case3012wp_k partitioned into twelve regions with 100 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 13: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case3012wp_k.
IV-G Case: pglib_opf_case4661_sdet
The pglib_opf_case4661_sdet dataset is a large-scale synthetic power system model with 4,661 buses. It was created using the Sustainable Data Evolution Technology tool and is publicly available through the eGRID Data portal. This model is part of the PGLib OPF v19.05 benchmark suite and is provided in the public domain to support research in scalable and realistic power system analysis.
The pglib_opf_case4661_sdet benchmark system was partitioned into fourteen electrically coherent regions using the proposed graph-based spectral clustering methodology. Before partitioning, the original bus numbering in the MATPOWER data was reordered to enable consistent and systematic region assignment. The underlying graph representation of the network was verified to be fully connected, ensuring that all components were eligible for inclusion in the partitioning process. After 10 iterations, the algorithm identified an optimal clustering configuration with 14 regions and 284 inter-regional tie-lines. The reference (slack) bus, originally located at bus 2625 in the MATPOWER data, was assigned to Region 3 (denoted as mpc_regionR3).
A validation step was performed after partitioning to confirm the structural integrity of the distributed dataset. It verified that all 4661 buses, 5997 branches, and 1176 generators from the original MATPOWER case were successfully preserved and assigned across the generated regional files. The partitioned data has been saved in both .mat and script-based .m formats under the name pglib_opf_case4661_sdet_14regions. A visual summary of the multi-regional topology is presented in Fig. 14. Subsequently, the distributed DC OPF and AC OPF solvers were executed on the partitioned system, and their convergence characteristics are shown through the worst primal residual plots in Figs. 15(a) and 15(b) with $\rho=4000$.
Figure 14: Topology of pglib_opf_case4661_sdet partitioned into fourteen regions with 284 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 15: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case4661_sdet.
IV-H Case: pglib_opf_case6468_rte
The pglib_opf_case6468_rte [20] dataset provides a realistic snapshot of the French high voltage and very high voltage transmission network in 2013. It contains 6,468 buses, 1,295 generators, and 9,000 branches, offering a highly detailed and complex test case for power system optimization. While the data are based on real grid topologies, they are intended solely for academic and research purposes and should not be used for operational or planning decisions related to the French or European power grids. This dataset is part of the IEEE PES Power Grid Library (PGLib-OPF v19.05) and was developed by Réseau de Transport d’Électricité (RTE), the French transmission system operator. It supports validating optimization algorithms, mathematical modeling techniques, and simulation tools in large-scale power system analysis.
The pglib_opf_case6468_rte benchmark was successfully partitioned into electrically coherent regions using our spectral clustering-based partitioning algorithm. This method ensured that the network’s graph structure remained fully connected. The best clustering outcome was identified within 10 iterations, resulting in 16 regions interconnected by 204 inter-regional tie-lines. Bus renumbering was applied to the original MATPOWER case to enable consistent indexing across the partitioning workflow. The system’s slack bus, labeled as bus 4733, was assigned to Region 6 (mpc_regionR6).
All buses, branches, and generators from the original MATPOWER dataset were accounted for in the partitioned output, confirming the completeness and validity of the partitioning. Specifically, the resulting distributed dataset includes all 6468 buses, 9000 branches, and 1295 generators from the original case. The partitioned version has been saved in both .mat and script-based .m formats under the name pglib_opf_case6468_rte_16regions. A visualization of the region-level partitioned topology is presented in Fig. 16. To validate the partition, we applied the distributed DC OPF and AC OPF solvers, which successfully converged. Their worst primal residual plots over ADMM iterations are shown in Fig. 17 with $\rho=6000$.
Figure 16: Topology of pglib_opf_case6468_rte partitioned into 16 regions with 204 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 17: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case6468_rte.
IV-I Case: pglib_opf_case9241_pegase
The pglib_opf_case9241_pegase dataset provides a highly detailed representation of the European high voltage transmission network. It includes 9,241 buses, 1,445 generators, and 16,049 branches operating across multiple voltage levels, including 750, 400, 380, 330, 220, 154, 150, 120, and 110 kV. Although the dataset is fictitious and does not reflect real-world data, it is valuable for validating optimization algorithms and simulation tools in large-scale power system research. The data were released as part of the Pan European Grid Advanced Simulation and State Estimation (PEGASE) project [21].
The pglib_opf_case9241_pegase benchmark represents one of the largest systems in the MATPOWER repository and was successfully partitioned using our graph-based spectral clustering methodology. The partitioning process ensured complete inclusion of all 9241 buses and confirmed the full connectivity of the underlying power grid graph. The best clustering configuration was determined within 10 iterations, resulting in 18 regions connected by 197 inter-regional tie-lines. The system’s reference bus, numbered 4231, was assigned to Region 17 (mpc_regionR17). The reorganization of the data preserves the integrity of the original MATPOWER structure, with a total of 16049 branches and 1445 generators fully accounted for across the distributed regions. The partitioned results were saved in both .mat and .m formats under the name pglib_opf_case9241_pegase_18regions. A visual representation of the multi-region topology is provided in Fig. 18. This dataset has been validated by executing the distributed DC OPF and AC OPF solvers, with the convergence of the worst primal residuals depicted in Fig. 19 with $\rho=8000$.
Figure 18: Topology of pglib_opf_case9241_pegase partitioned into 18 regions with 197 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 19: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case9241_pegase.
V Conclusion
This work presented DPLib, a reproducible and open-source benchmark library tailored for distributed and decentralized power system studies. DPLib addresses the gap in standardized datasets for distributed OPF research by offering a graph-based partitioning toolkit and over 20 multi-region test cases derived from MATPOWER. The library is complemented by modular ADMM-based DC and AC OPF solvers, which enable scalable validation across diverse case sizes. Numerical results demonstrate the accuracy of the generated test cases and convergence of the solvers, making DPLib a practical foundation for future research in scalable, privacy-preserving, and resilient power system analysis.

### v3

@@
 Distributed and decentralized methods offer scalability, privacy preservation, and resilience to single points of failure, making them increasingly important for modern power systems.
-However, unlike centralized tools such as MATPOWER, no general-purpose, reproducible data library package currently exists for distributed power system studies.
-DPLib, available at GitHub, fills this gap by providing a standard power system library featuring over 20 multi-region benchmark test cases of varying sizes, along with a graph-based partitioning toolkit that decomposes any MATPOWER test system into multiple electrically coherent regions.
+However, unlike centralized tools such as PGLib, no general-purpose, reproducible data library package currently exists for distributed power system studies.
+DPLib, available at GitHub, fills this gap by providing a standard power system library featuring over 20 multi-region benchmark test cases of varying sizes, along with a graph-based partitioning toolkit that decomposes any MATPOWER test system into multiple regions.
 The partitioning toolkit, an easy-to-use MATLAB code, generates standardized .mat and .m files, along with region visualizations for intuitive understanding.
-We also provide modular, easy-to-use distributed optimal power flow (OPF) solvers: an alternating direction method of multipliers(ADMM)-based DC-OPF solver implemented in YALMIP, and an ADMM-based AC-OPF solver leveraging IPOPT.
+We also provide modular, easy-to-use distributed optimal power flow (OPF) solvers: an alternating direction method of multipliers (ADMM)-based DC OPF solver implemented in YALMIP, and an ADMM-based AC OPF solver leveraging IPOPT.
 These solvers validate the generated test systems for distributed optimization applications.
 Numerical results validate the generated test cases, establishing DPLib as a foundation for reproducible distributed power system research.
-Index Terms:
-Distributed power grid test systems, standard benchmark library, MATPOWER, distributed optimization.
 I Introduction
-Centralized frameworks have long served as the dominant approach for modeling and solving large-scale power system problems, including optimal power flow (OPF) and security-constrained unit commitment.
+Centralized frameworks have long served as the dominant approach for modeling and solving large-scale power system problems, including optimal power flow (OPF), planning, control, and unit commitment.
 A key enabler behind the success and widespread adoption of centralized optimization and analysis in power systems is the availability of standardized benchmark test cases and libraries.
-Among these, MATPOWER has emerged as a widely accepted benchmark tool, offering consistent, well-structured data for centralized OPF studies  [1].
-Its role in enabling reproducibility, fair performance comparisons, and methodological development has been fundamental to the evolution of centralized power system optimization and analysis.
-For centralized OPF studies, for instance, [2] proposed convex relaxations for AC OPF and validated their results on MATPOWER test cases.
+Among these, MATPOWER has emerged as a widely accepted benchmark tool, offering consistent, well-structured data for centralized studies [22].
+Its role in enabling reproducibility, fair performance comparisons, and methodological development has been fundamental to the evolution of centralized power system analysis.
+For centralized studies, for instance, [11] proposed convex relaxations for AC OPF and validated their results on MATPOWER test cases.
 [16] provided a comprehensive review of relaxation techniques for OPF, with all benchmarks sourced from MATPOWER.
@@
 In the last decade, distributed and decentralized computation has gained significant momentum in the power systems community, especially in applications such as distributed OPF, voltage control, state estimation, and demand-side management [15, 10, 6].
-These include studies that apply alternating direction method of multipliers (ADMM) decomposition [8, 10], distributed voltage control [11], parallel and distributed OPF [12, 13], consensus-based optimization [14], and cooperative control frameworks [15, 16].
+These include studies that apply alternating direction method of multipliers (ADMM) decomposition [10, 4], distributed voltage control [14], parallel and distributed OPF [21, 8], consensus-based optimization [18], and cooperative control frameworks [3, 7, 9].
 Additional contributions include scalable distributed frameworks for transmission networks [12], graph-based coordination strategies [15], and distributed control in microgrids [5].
@@
 In response to this need, we develop a distributed power system library, DPLib, that addresses the lack of standardized resources for distributed power system studies by providing a comprehensive library of over 20 benchmark test cases, ranging from 73 to 9241 buses.
-These multi-region systems are generated using the DPLib graph-based partitioning toolkit that ensures electrical coherence and captures the corresponding multi-regional network topology.
+These multi-region systems are generated using the DPLib graph-based partitioning toolkit that captures the corresponding multi-regional network topology.
 Each case includes the necessary structural and connectivity data to support distributed analysis.
 All test cases have been verified using distributed OPF problems to demonstrate their practicality and confirm their suitability for reproducible and scalable distributed optimization.
-In addition to a standard library, a key requirement is a general and user-friendly partitioning tool capable of processing any standard MATPOWER case and dividing it into electrically coherent subregions based on user-defined parameters, such as the number of regions.
+In addition to a standard library, a key requirement is a general and user-friendly partitioning tool capable of processing any standard MATPOWER case and dividing it into subregions based on user-defined parameters, such as the number of regions.
 An effective tool should preserve the integrity of the original network, identify and document inter-regional tie-lines, and produce output in a structured format consistent with centralized MATPOWER cases.
@@
 The DPLib graph-based partitioning toolkit fulfills all of these requirements.
-While the distributed library includes several pre-partitioned and validated datasets, researchers may also wish to generate and test their customized partitions using the DPLib graph-based toolkit.
+While the distributed library includes several pre-partitioned and validated datasets, researchers may also wish to test their customized partitions generated by the DPLib graph-based toolkit.
 To support this, DPLib provides ADMM-based solvers for both DC OPF and AC OPF that are fully compatible with the partitioned data structure produced by the tool.
@@
 The main contribution of this work is the introduction of DPLib, a unified framework that establishes the first standard distributed data library for power system applications.
-Besides the library, DPLib consists of three core tools: a graph-based partitioning tool that decomposes any MATPOWER case into multiple regions using spectral clustering and k-means, enabling the creation of distributed datasets; a modular ADMM-based distributed DC OPF solver implemented in YALMIP; and a distributed AC OPF solver using ADMM and IPOPT.
+Besides the library, DPLib consists of three core tools: a graph-based partitioning tool that decomposes any MATPOWER case into multiple regions using spectral clustering and k-means, enabling the creation of distributed datasets; a modular ADMM-based distributed DC OPF solver implemented in YALMIP [13]; and a distributed AC OPF solver using ADMM and IPOPT [20].
 The DC OPF and AC OPF solvers are designed to validate and benchmark the partitioned datasets under realistic operating conditions.
@@
 II-A Power Network Modeling and Spectral Clustering
-To enable distributed optimization and control, a centralized network must first be divided into smaller regions that remain physically meaningful and electrically coherent.
-This partitioning process starts by modeling the power system as an undirected graph.
-We represent the power network as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where:
-  • $\mathcal{V}$ is the set of buses (nodes), as defined in the original MATPOWER case file.
-  • $\mathcal{E}$ is the set of transmission lines (edges), each connecting two buses.
-The graph is undirected because the power flow direction does not impact the structure for partitioning purposes—only a connection between buses matters.
-A path in a graph is a sequence of nodes such that an edge connects each consecutive pair of nodes in the sequence.
-A connected graph is an undirected graph with a path between every pair of nodes.
-This means it is possible to travel from any node to any other by following a sequence of edges in the graph.
-To describe the structure of $\mathcal{G}$, we construct the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $n=|\mathcal{V}|$ is the total number of buses.
-The entries of this matrix are defined as:
-$A_{ij}=\begin{cases}1,&\text{if there is a line connecting bus }i\text{ and bus }j,\\ 0,&\text{otherwise}.\end{cases}$
-Because the graph is undirected, $A$ is symmetric.
-Next, we define the degree matrix $D\in\mathbb{R}^{n\times n}$, a diagonal matrix where each diagonal entry $D_{ii}$ equals the number of lines connected to bus $i$:
-$D_{ii}=\sum_{j=1}^{n}A_{ij}.$
-Using these two matrices, we define the graph Laplacian $L=D-A.$ The Laplacian matrix $L$ encodes the power network connectivity and topology.
+To prepare the network for spectral partitioning, the power system is modeled as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where buses form the nodes and transmission lines form the edges.
+The graph representation is encoded through the adjacency matrix $A$, and two forms of this matrix may be constructed depending on how connectivity is interpreted.
+In its simplest form, the network is viewed purely through its topological structure. Each transmission line indicates the presence of a connection, and the adjacency matrix is as
+$A_{ij}=\begin{cases}1,&\text{if a line connects buses }i\text{ and }j,\\ 0,&\text{otherwise}.\end{cases}$ (1)
+This unweighted representation treats all lines uniformly and emphasizes the combinatorial shape of the network.
+Such a formulation is robust to data irregularities, reflects the way regional boundaries appear in practice when tie-lines are sparse, and preserves the graph-theoretic properties that spectral clustering exploits most directly.
+A second option introduces electrical heterogeneity explicitly.
+In this case, the weight of an edge is determined by parameters such as line reactance $x_{ij}$, line charging susceptance $b_{ij}$, and thermal rating $S_{\max,ij}$.
+After appropriate normalization, a composite quantity is assigned to each existing edge,
+$A_{ij}=w(x_{ij},b_{ij},S_{\max,ij}),$ (2)
+so the graph captures not only the presence of a connection but also its relative electrical strength.
+This weighted representation highlights tightly coupled corridors and down-weights weak or limited-capacity lines, and therefore encodes additional physical detail into the graph used for clustering.
+Although the toolkit supports a weighted Laplacian that incorporates electrical heterogeneity, we treat this option as an auxiliary feature rather than the default mode.
+Weighted formulations can highlight electrically strong corridors, but doing so requires a choice of normalization, scaling, and aggregation of heterogeneous parameters—decisions that are not standardized across power-system studies.
+Because the purpose of DPLib is to produce baseline multi-region datasets for benchmarking distributed algorithms, we use the unweighted formulation to ensure that all results are comparable, repeatable, and free of case-dependent parameter tuning.
+Researchers who wish to emphasize electrical strength rather than topological connectivity may enable the weighted variant, which the toolkit provides for completeness.
+Both constructions lead to the degree matrix $D$ defined by $D_{ii}=\sum_{j}A_{ij}$, and the Laplacian L=DA.L=D-A.
+The Laplacian matrix $L$ encodes the power network connectivity and topology.
 For a connected graph, $L$ is symmetric and positive semi-definite, meaning all its eigenvalues are non-negative.
 Importantly, the smallest eigenvalue of $L$ is always zero, corresponding to the trivial uniform eigenvector.
-To divide the system into $k$ electrically coherent regions, we use spectral clustering, which leverages the properties of $L$.
+To divide the system into $k$ regions, we use spectral clustering, which leverages the properties of $L$.
 The first step is to compute the eigenvalues and eigenvectors of $L$.
@@
 The clustering result yields a partition of all buses into $k$ regions.
-Each cluster groups strongly connected buses, leading to electrically coherent and balanced regions.
+Each cluster groups strongly connected buses, leading to balanced regions where each has its generation.
 After forming the clusters, inter-regional tie-lines (lines that connect buses belonging to different clusters) are identified.
-To improve quality, the algorithm may repeat the clustering multiple times and select the configuration that minimizes the number of inter-regional tie-lines.
-This spectral clustering approach provides a systematic and scalable method for decomposing any MATPOWER case into meaningful regions.
-It aligns well with practical expectations since most real-world partitioned power networks are formed based on geographic or electrical proximity.
-A tool that can automatically convert MATPOWER test cases into $k$ well-structured regions is essential for distributed power system studies.
-The DPLib partitioning toolkit fills this gap by generating standardized multi-region datasets from any MATPOWER case, supporting reproducibility and scalability.
+To improve quality, the algorithm repeats the clustering multiple times and select the configuration that minimizes the number of inter-regional tie-lines.
 II-B DPLib Partitioning Toolkit
-The DPLib graph-based partitioning toolkit implements the method described above and automates the generation of partitioned MATPOWER cases.
-The partitioning toolkit is publicly available at GitHub, including the MATLAB implementation file named partitioning_code.m.
-The code uses MATLAB R2020a or newer, along with the MATPOWER toolbox.
-Upon running the script, the user is prompted to enter the name of a valid MATPOWER case file (for example, case118) and specify the desired number of regions into which the system should be divided.
-Internally, the code loads the MATPOWER case and checks whether the bus numbers are sequential.
-If not, it renumbers the buses from one to $n$ and updates all corresponding references in the branch and generator matrices.
-It then constructs the undirected adjacency matrix based on the network’s branch connectivity and forms the corresponding graph Laplacian.
-The eigenvectors of the Laplacian are computed and used as features for $k$-means clustering.
-The clustering process is repeated multiple times to ensure optimality, and the configuration with the lowest number of inter-regional tie-lines is selected.
-For each candidate partitioning, regional subcases are formed by extracting internal buses, branches, generators, and generation cost data for each region.
-The script ensures that each region has at least one generator.
-Buses are then renumbered locally within each region to maintain consistency.
-Inter-regional tie-lines are identified as any branch connecting buses in different regions and stored.
-Once a feasible partitioning is found, the code saves the result in multiple formats.
-A MATLAB .mat file is created, containing all regional subcases under separate variables (e.g., mpc_regionR1, mpc_regionR2, etc.), along with a comprehensive table of all inter-regional tie-lines labeled by their originating and receiving regions and complete branch data.
-A corresponding .m file is also generated, exporting the same structures in a human-readable script format suitable for inspection or direct use in other codes.
-Additionally, the code creates a graphical visualization of the partitioned system, where each region is displayed as a labeled node in a circular layout, and tie-lines are represented as edges connecting the areas.
-The number of tie-lines between each pair of regions is also annotated.
-During execution, the code prints detailed messages to the MATLAB command window.
-These include whether bus renumbering was necessary, confirmation of full graph connectivity, the number of generated regions and tie-lines, the region assignment of the reference (slack) bus, and whether all buses, branches, and generators are fully included in the partitioned result.
-It also confirms the successful saving of the generated .mat, .m, and .png files.
-The algorithm of the partitioning toolkit, illustrating the complete data processing pipeline, is provided in 1.
+The DPLib partitioning toolkit implements the graph-based clustering framework described in the previous subsection and automates the full workflow for constructing multi-region MATPOWER datasets.
+The toolkit supports both adjacency representations introduced earlier, allowing the Laplacian to be formed either from the purely topological structure of the network or from a weighted graph that incorporates electrical parameters.
+In the unweighted mode, each transmission line contributes a unit entry to the adjacency matrix, so the Laplacian reflects the combinatorial connectivity of the system.
+In the weighted mode, each existing edge is assigned a composite weight constructed from line reactance, line charging susceptance, and thermal capacity after normalization, allowing the Laplacian to embed relative electrical strength directly into the spectral geometry.
+Although both modes are supported, the unweighted Laplacian is used as the default for generating standardized multi-region test cases.
+This choice ensures that the resulting partitions depend only on the physical topology of the MATPOWER system and remain independent of parameter scaling, model variations, or dataset-dependent numerical ranges.
+The unweighted construction also aligns with how regional boundaries appear in real-world transmission networks, where geographical separations and sparse tie-lines primarily determine structural decomposition.
+Using the unweighted version therefore provides consistent, reproducible partitions across different datasets and preserves compatibility with studies that compare methods on topologically equivalent cases.
+The workflow proceeds as follows.
+The MATPOWER case is loaded, and if the original bus numbering is not sequential, all buses, branches, and generators are renumbered from $1$ to $n$.
+Depending on the selected mode, the adjacency matrix is constructed using either the binary definition or the weighted formulation.
+The degree matrix and Laplacian are then formed, and the $k$ Laplacian eigenvectors associated with the smallest nonzero eigenvalues are computed to obtain the spectral embedding.
+Multiple runs of $k$-means are applied to the rows of this matrix to obtain a stable clustering, and the configuration with the smallest number of inter-regional tie-lines is selected.
+For the chosen partition, regional subcases are created by extracting internal buses, branches, generators, and cost data, and each region is locally renumbered to maintain consistency.
+Tie-lines connecting different regions are identified and stored together with their full MATPOWER branch data.
+The completed partition is exported in several formats.
+A .mat file contains all regional subcases and a table describing every inter-regional tie-line.
+A corresponding .m script reproduces the same structures in a clear and readable format.
+A graphical visualization is produced in .png format, showing regions arranged in a force-directed layout together with annotated tie-lines.
+These outputs provide a consistent, reproducible, and easily verifiable multi-region dataset that can be used directly in distributed studies.
+The algorithm of the partitioning toolkit, illustrating the complete data processing pipeline, is provided in Algorithm 1.
 Algorithm 1 DPLib Partitioning Toolkit
-1:User Input:
-  • Name of the original MATPOWER case (e.g., case118)
-  • Number of desired regions $k$
-2: Load the MATPOWER case file
-3: Check bus numbering; if non-sequential, renumber from 1 to $n$ and update all references
-4: Construct adjacency matrix $A$ from the branch data.
-5: Construct degree matrix $D$ with $D_{ii}=\sum_{j}A_{ij}$
-6: Form graph Laplacian $L=D-A$
-7: Compute the first $k$ nonzero eigenvectors of $L$
-8: Stack eigenvectors into matrix $U\in\mathbb{R}^{n\times k}$
-9: Apply $k$-means clustering to the rows of $U$
-10: Repeat clustering for multiple iterations to minimize the number of tie-lines
-11:for each region (cluster) do
-12:  Extract internal buses, branches, generators, and cost data
-13:  Ensure that at least one generator exists in each region
-14:  Locally renumber buses
-15:  Identify inter-regional tie-lines
-16:end for
-17: Save output in the following formats:
-  • .mat file with region-wise MATPOWER data
-  • .m script file exporting the same data
-  • .png visualization showing regions and tie-lines
+1: Input MATPOWER case name and desired number of regions $k$
+2: Load MATPOWER data; if bus numbering is non-sequential, renumber buses and update references
+3: Construct adjacency matrix AA:
+  1. 1.
+    Unweighted: $A_{ij}=1$ if a branch connects $i$ and $j$
+  2. 2.
+    Weighted: assign $A_{ij}=w(x_{ij},b_{ij},S_{\max,ij})$
+4: Form degree matrix $D$ and Laplacian L=DAL=D-A
+5: Compute the Laplacian eigenvectors associated with the $k$ smallest nonzero eigenvalues and form $U$
+6: For multiple attempts: apply $k$-means to rows of $U$, build candidate regions, identify tie-lines, and discard candidates without generators
+7: Select the valid partition with the minimum number of tie-lines
+8:for each region do
+9:  Extract internal buses, branches, generators, and cost data; locally renumber buses
+10:end for
+11: Identify the region of the slack bus and record all inter-regional tie-lines
+12: Export regional cases to .mat and .m files and generate topology visualization
 III Validating DPLib Test Systems Using OPF
-To ensure that the generated multi-region test cases are electrically coherent and suitable for distributed power system studies, it is essential to validate them through power system analysis.
-Although the DPLib partitioning toolkit generates distributed variants of MATPOWER test systems applicable to a wide range of power system studies, including control, state estimation, or stability analysis, this section evaluates these test systems for distributed OPF applications.
+To ensure that the generated multi-region test cases are suitable for distributed power system studies, it is essential to validate them through power system analysis.
+Although the DPLib partitioning toolkit generates distributed variants of MATPOWER test systems applicable to a wide range of power system studies, including control, state estimation, or stability analysis, this section evaluates these test systems for distributed OPF applications only.
 OPF has been selected as a representative benchmarking problem because of its maturity, availability of centralized references, and compatibility with distributed solvers due to over a decade of research in this area.
@@
 We formulate the OPF problem in both DC and AC settings for each region $r$.
-Each region solves a local OPF problem while maintaining consistency at the boundaries via duplicated (virtual) variables and tie-line constraints.
+Each region solves a local OPF problem while maintaining consistency at the boundaries via duplicated (virtual) variables and tie-line constraints [8].
 These virtual variables (buses) enable each region to model tie-line interactions locally without requiring direct enforcement of global constraints.
@@
 III-A1 DC OPF Formulation
-The DC OPF for region $r$ is:
+The DC OPF for region $r$ is [8]:
 $\min_{\boldsymbol{x}^{r}}\sum_{i\in\Omega^{r}_{g}}f(p^{g}_{i})$ (3a)
@@
 $\qquad\quad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\frac{\theta_{i}-\theta_{j}}{x_{ij}}+\sum_{j\in\Omega^{r,\textrm{in}}_{i}}\frac{\theta_{i}-\tilde{\theta}_{j}}{x_{ij}},\quad\forall i\in\Omega^{r}_{b}$ (3c)
-$\qquad\quad\underline{p}_{ij}\leqslant p_{ij}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (1d)
-$\qquad\quad\theta_{\text{ref }}=0,$ (1e)
-$\qquad\quad\theta_{i}=\tilde{\theta}_{i},\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (1f)
-$\boldsymbol{x}^{r}=[\boldsymbol{p}^{g\top},\boldsymbol{\theta}^{\top}]\in\mathbb{R}^{n^{r}_{g}+n^{r}_{b}}$ is the decision vector for region $r$, including generation $p^{g}_{i}$ and voltage angles $\theta_{i}$.
-Equation (1a) minimizes the local cost $f(p^{g}_{i})$.
-Constraints (1b)–(1c) enforce generator limits and power balance using demand $p^{d}_{i}$, line reactance $x_{ij}$, and virtual angles $\tilde{\theta}_{j}$.
-Constraint (1d) bounds internal and tie-line flows $p_{ij}$, and (1f) ensures consistency between real and virtual bus angles.
+$\qquad\quad\underline{p}_{ij}\leqslant\frac{\theta_{i}-\theta_{j}}{x_{ij}}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r}_{l}$ (3d)
+$\qquad\quad\underline{p}_{ij}\leqslant\frac{\theta_{i}-\tilde{\theta}_{j}}{x_{ij}}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (3e)
+$\qquad\quad\theta_{\text{ref }}=0,$ (3f)
+$\qquad\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (3g)
+$\qquad\quad\theta_{i}=\tilde{\theta}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (3h)
+The decision vector for region r contains the active power generations $p^{g}_{i}$ for all generating units $i\in\Omega^{r}_{g}$ and the bus voltage phase angles $\theta_{i}$ for all buses $i\in\Omega^{r}_{b}$, giving $x^{r}\in\mathbb{R}^{n^{r}_{g}+n^{r}_{b}}$.
+The objective in (3a) minimizes the local generation cost $f(p^{g}_{i})$.
+Constraint (3b) applies the minimum and maximum generation limits for all units in the set $\Omega^{r}_{g}$.
+The nodal balance equation (3c) uses the demand $p^{d}_{i}$, the reactance $x_{ij}$ of each line, the set of internal neighboring buses $\Omega^{r}_{i}$, and the set of inter-regional neighboring buses $\Omega^{r,\mathrm{in}}_{i}$, and incorporates virtual boundary angles $\tilde{\theta}_{j}$ that represent buses belonging to adjacent regions.
+Constraints (3d)–(3e) impose the power-flow limits $\underline{p}_{ij}$ and $\bar{p}_{ij}$ on all internal lines in $\Omega^{r}_{l}$ and all tie-lines in $\Omega^{r,\mathrm{in}}_{l}$.
+Constraint (3f) selects the regional reference bus.
+Finally, constraints (3g)–(3h) enforce angle consistency across inter-regional connections by matching the real boundary-bus angles with their corresponding virtual angles.
 III-A2 AC OPF Formulation
-The AC OPF for region $r$ is:
+The AC OPF for region $r$ [8]:
 $\min_{\boldsymbol{x}^{r}}\sum_{i\in\Omega^{r}_{g}}f(p^{g}_{i})$ (4a)
 s.t.
-$\qquad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\underline{q}_{i}^{g}\leqslant q^{g}_{i}\leqslant\bar{q}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (2b)
-$\qquad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(p_{ij,f}+p_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}\tilde{p}_{ij},\quad\forall i\in\Omega^{r}_{b}$ (2c)
-$\qquad q^{g}_{i}-q^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(q_{ij,f}+q_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}\tilde{q}_{ij},\quad\forall i\in\Omega^{r}_{b}$ (2d)
+$\qquad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (4b)
+$\qquad\underline{q}_{i}^{g}\leqslant q^{g}_{i}\leqslant\bar{q}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (4c)
+$\qquad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(p_{ij,f}+p_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}{p}_{ij,\xi},\quad\forall i\in\Omega^{r}_{b}$ (4d)
+$\qquad q^{g}_{i}-q^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(q_{ij,f}+q_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}{q}_{ij,\xi},\quad\forall i\in\Omega^{r}_{b}$ (4e)
 $\qquad p_{ij,f}=G_{ij,1}v_{i}^{2}+G_{ij,2}v_{i}v_{j}\cos(\theta_{i}-\theta_{j})$
-$\qquad\qquad\quad+B_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (2e)
+$\qquad\qquad\quad+B_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (4f)
 $\qquad q_{ij,f}=-B_{ij,1}v_{i}^{2}-B_{ij,2}v_{i}v_{j}\cos(\theta_{i}-\theta_{j})$
-$\qquad\qquad\quad+G_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (2f)
+$\qquad\qquad\quad+G_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (4g)
 $\qquad p_{ij,t}=G_{ij,4}v_{j}^{2}+G_{ij,3}v_{i}v_{j}\cos(\theta_{j}-\theta_{i})$
-$\qquad\qquad\quad+B_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (2g)
+$\qquad\qquad\quad+B_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (4h)
 $\qquad q_{ij,t}=-B_{ij,4}v_{j}^{2}-B_{ij,3}v_{i}v_{j}\cos(\theta_{j}-\theta_{i})$
-$\qquad\qquad\quad+G_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (2h)
-$\qquad p_{ij,f}^{2}+q_{ij,f}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2i)
-$\qquad p_{ij,t}^{2}+q_{ij,t}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2j)
-$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\theta_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2k)
-$\qquad\theta_{\text{ref}}=0$ (2l)
-$\qquad\theta_{i}=\tilde{\theta}_{i},\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (2m)
-$\qquad v_{i}=\tilde{v}_{i},\quad\tilde{v}_{j}=v_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (2n)
-The decision vector is $\boldsymbol{x}^{r}=[\boldsymbol{p}^{g\top},\boldsymbol{q}^{g\top},\boldsymbol{\theta}^{\top},\boldsymbol{v}^{\top}]$, including generation $(p^{g}_{i},q^{g}_{i})$, angles $\theta_{i}$, and voltages $v_{i}$ at region $r$.
-Constraints (2b)–(2d) enforce generator and nodal power limits; (2e)–(2h) define AC flows using admittance terms $G_{ij,k},B_{ij,k}$.
-Limits on flows and angle differences are imposed in (2i)–(2k).
-Reference angle (2l) and virtual bus equalities (2m)–(2n) ensure boundary consistency.
-III-A3 ADMM Coordination
-To ensure consistency between neighboring regions, we use ADMM to coordinate virtual variables [10]. For example, the voltage angle consistency is enforced via:
-$L_{\rho}(\theta,\tilde{\theta},\lambda)=f(p^{g})+\lambda^{\top}(\theta-\tilde{\theta})+\frac{\rho}{2}\|\theta-\tilde{\theta}\|^{2},$ (3)
-where $\lambda$ are the dual variables and $\rho$ is the penalty parameter. Similar terms are added for voltage magnitudes in AC OPF.
-Each ADMM iteration proceeds as:
-  1. 1.
-    Solve each region’s local OPF using the current dual variables.
-  2. 2.
-    Update dual variables based on tie-line consistency.
-  3. 3.
-    Check primal residuals and stop if below a threshold.
-In general, coordination across regions is achieved using the ADMM.
-The distributed DC OPF and AC OPF formulations in (1) and (2) incorporate local objectives and tie-line consistency via augmented Lagrangian.
-Subproblems are solved locally at each iteration, dual variables are updated, and convergence is checked via the worst primal residual.
+$\qquad\qquad\quad+G_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (4i)
+$\qquad p_{ij,f}^{2}+q_{ij,f}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}$ (4j)
+$\qquad p_{ij,t}^{2}+q_{ij,t}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}$ (4k)
+$\qquad{p}^{2}_{ij,\xi}+{q}^{2}_{ij,\xi}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (4l)
+$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\theta_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r}_{l}$ (4m)
+$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\tilde{\theta}_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (4n)
+$\qquad\theta_{\text{ref}}=0,$ (4o)
+$\qquad\tilde{\theta}_{j}=\theta_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (4p)
+$\qquad\theta_{i}=\tilde{\theta}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (4q)
+$\qquad\tilde{v}_{j}=v_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (4r)
+$\qquad v_{i}=\tilde{v}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (4s)
+The AC OPF subproblem for region r uses the same generation set $\Omega^{r}_{g}$ and bus set $\Omega^{r}_{b}$, where the decision variables include active and reactive generations $p^{g}_{i}$ and $q^{g}_{i}$, bus voltage magnitudes $v_{i}$, and phase angles $\theta_{i}$, together with virtual boundary variables $\tilde{\theta}_{j}$ and $\tilde{v}_{j}$ for all inter-regional neighbors in $\Omega^{r,\mathrm{in}}_{i}$.
+The objective in (4a) minimizes the generation cost over all units in $\Omega^{r}_{g}$.
+Generator active and reactive limits are enforced in (4b)–(4c).
+Power-balance constraints (4d)–(4e) use local demands $p^{d}_{i}$ and $q^{d}_{i}$, as well as flows from internal neighbors $j\in\Omega^{r}_{i}$ and inter-regional neighbors $j\in\Omega^{r,\mathrm{in}}_{i}$, where $p_{ij,f},p_{ij,t},q_{ij,f},q_{ij,t}$ represent the AC branch flows defined by the conductance and susceptance parameters $G_{ij,k}$ and $B_{ij,k}$ in (4f)–(4i).
+Thermal limits for internal lines $\Omega^{r}_{l}$ and tie-lines $\Omega^{r,\mathrm{in}}_{l}$ are enforced in (4j)–(4l).
+The phase-angle difference bounds in (4m)–(4n) use the parameters $\underline{\Delta\theta}_{ij}$ and $\overline{\Delta\theta}_{ij}$, and (4o) sets the reference bus angle.
+Finally, constraints (4p)–(4s) ensure consistency between the real and virtual boundary-bus variables, matching $\theta_{i}$ with $\tilde{\theta}_{i}$ and $v_{i}$ with $\tilde{v}_{i}$ across all inter-regional connections in $\Omega^{r,\mathrm{in}}_{l}$.
+III-A3 ADMM Decomposition
+The distributed OPF solver uses the sequential form of ADMM, where regions are updated one after another in a Gauss–Seidel manner.
+Each region solves its local OPF problem using the most recent boundary voltage information from neighbors that have already been updated in the current iteration, while using the previous-iteration values for neighbors that have not yet been updated.
+The shared variables are the voltage phase angles and voltage magnitudes of boundary buses.
+Their virtual copies represent the neighboring regions’ views of these boundary variables.
+Before the ADMM iterations start, we also compute problem-dependent scaling factors for the boundary mismatches.
+For the phase angles, a characteristic scale $\theta_{\text{scale}}$ is obtained from the line reactances and thermal limits as
+$\theta_{\max,\ell}\approx|x_{\ell}|\,\frac{\text{rateA}_{\ell}}{\text{baseMVA}},$
+and $\theta_{\text{scale}}$ is chosen as a robust statistic (for example, the median) of $\{\theta_{\max,\ell}\}$ over all lines.
+For AC OPF, an additional voltage-magnitude scale $v_{\text{scale}}$ is defined from the bus voltage limits, for example as half the spread between the minimum and maximum allowed magnitudes.
+These scales are used to normalize the consensus residuals and make them dimensionless.
+Consistency across regions is enforced by an augmented Lagrangian of the form
+$L_{\rho}(\theta,\tilde{\theta},v,\tilde{v},\lambda_{\theta},\lambda_{v})=$ $f(p^{g})+\lambda_{\theta}^{\top}(\frac{\theta-\tilde{\theta}}{\theta_{\text{scale}}})+\lambda_{v}^{\top}(\frac{v-\tilde{v}}{v_{\text{scale}}})$
+$+\tfrac{\rho_{\theta}}{2}\|\frac{\theta-\tilde{\theta}}{\theta_{\text{scale}}}\|^{2}+\tfrac{\rho_{v}}{2}\|\frac{v-\tilde{v}}{v_{\text{scale}}}\|^{2},$ (5)
+[…]
+During iteration $k$, region $r$ performs the primal update by solving its local AC or DC OPF subproblem.
+The virtual variables from neighbors are fixed parameters in this optimization.
+After the OPF solve, region $r$ updates its real boundary voltages $(\theta_{r}^{k+1},v_{r}^{k+1})$ and refreshes its virtual copies $(\tilde{\theta}_{r}^{k+1},\tilde{v}_{r}^{k+1})$ using the newest available values from neighboring regions.
+This constitutes the primal update step of ADMM.
+Once all regions finish their primal updates, the dual variables are updated based on normalized residuals.
+For each tie-line boundary bus, the updates are
+$\lambda_{\theta}^{k+1}$ $=\lambda_{\theta}^{k}+\rho_{\theta}\frac{\theta^{k+1}-\tilde{\theta}^{k+1}}{\theta_{\text{scale}}},$ (6)
+$\lambda_{v}^{k+1}$ $=\lambda_{v}^{k}+\rho_{v}\frac{v^{k+1}-\tilde{v}^{k+1}}{v_{\text{scale}}}.$ (7)
+In the DC-OPF case only the angle channel is present, so the update for $\lambda_{v}$ and the associated terms are omitted.
+Convergence is assessed using the worst-case normalized primal residual across all tie-lines. The primal residual at iteration $k+1$ is computed as
+$r_{\text{pri}}^{k+1}=\max_{(i,j)\in\mathcal{E}_{\text{tie}}}\left\{\frac{|\theta_{i}^{k+1}-\tilde{\theta}_{i}^{k+1}|}{\theta_{\text{scale}}},\,\frac{|v_{i}^{k+1}-\tilde{v}_{i}^{k+1}|}{v_{\text{scale}}}\right\},$
+matching the residual definition used in [8], up to the normalization by $(\theta_{\text{scale}},V_{\text{scale}})$.
+In the DC-OPF case only the angle term is used in the maximum.
+The ADMM iterations terminate once this residual falls below a prescribed tolerance.
+To improve robustness across different network topologies and operating points, the penalty parameters are updated adaptively based on the relative sizes of the primal and dual residuals, following the heuristic in [17].
+Let $r_{\text{pri}}^{k}$ and $r_{\text{dual}}^{k}$ denote the normalized primal and dual residuals at iteration $k$.
+For the distributed DC-OPF solver, a single penalty parameter is updated according to
+$\rho^{k+1}=\begin{cases}\tau_{\text{incr}}\,\rho^{k},&\text{if }r_{\text{pri}}^{k}>\mu\,r_{\text{dual}}^{k},\\[1.99997pt] \rho^{k}/\tau_{\text{decr}},&\text{if }r_{\text{dual}}^{k}>\mu\,r_{\text{pri}}^{k},\\[1.99997pt] \rho^{k},&\text{otherwise},\end{cases}$
+with chosen parameters $\mu$, $\tau_{\text{incr}}$, and $\tau_{\text{decr}}$.
+For the distributed AC-OPF solver, the same heuristic is applied separately to the angle and magnitude channels, i.e., to $\rho_{\theta}$ and $\rho_{v}$, using different balance factors for the $\theta$ and $v$ residuals.
+To prevent numerical blow-up and avoid instability caused by excessively large penalties, we enforce an upper bound on the adaptive updates.
+Specifically, the penalty updates are terminated once the penalty reaches a prescribed maximum value: $\rho_{\mathrm{DC}}^{\max}$ for the distributed DC-OPF solver and $\rho_{\mathrm{AC}}^{\max}$ for the distributed AC-OPF solver.
+In practice, if an update would increase the current penalty beyond the corresponding cap, we set $\rho\leftarrow\rho_{\mathrm{DC}}^{\max}$ (DC) or $\rho_{\theta},\rho_{v}\leftarrow\rho_{\mathrm{AC}}^{\max}$ (AC) and disable further penalty adaptation thereafter.
+This safeguard ensures the algorithm does not diverge due to runaway growth of $\rho$.
+Moreover, to avoid overly frequent parameter changes—which can introduce oscillations and hinder convergence—we apply a cooldown period after each penalty update.
+Concretely, once $\rho$ (DC) or $\rho_{\theta},\rho_{v}$ (AC) is updated, we postpone any subsequent residual-balance checks and penalty updates for the next $N_{\mathrm{wait}}$ iterations, where $N_{\mathrm{wait}}=20$ for cases with fewer than $2000$ buses and $N_{\mathrm{wait}}=40$ for cases with $2000$ buses or more.
+This heuristic empirically stabilizes the convergence behavior by allowing the primal and dual iterates to settle under the new penalty before further adaptation is attempted.
+Overall, this adaptive strategy automatically increases the penalty when the primal residual dominates and decreases it when the dual residual dominates, promoting a better balance between consensus enforcement and dual convergence while remaining numerically stable through the maximum-penalty cap and the update cooldown mechanism.
+Figure 2 summarizes the sequential ADMM workflow, highlighting the order of local OPF solves, voltage exchange, dual updates (with scaled and adaptively tuned penalties), and residual checks.
+Initialize ADMMFor $r=1,\dots,K$: solve regional OPF and update boundary voltagesUpdate dual variables on tie-lines (scaled residuals, adaptive $\rho$)Compute worst-case normalized primal residualConverged?StopNext ADMM iterationyesno
+Figure 2: Sequential ADMM loop for distributed OPF with scaled residuals and adaptive penalties
 III-B General ADMM-Based OPF Solvers
@@
 III-B1 DPLib DC OPF Solver
-The DPLib DC OPF solves the OPF problem using ADMM by optimizing each region independently and enforcing agreement on shared variables (tie-line voltage angles) through iterative dual updates.
-The solver is available at GitHub, including the MATLAB implementation file named DistDCOPF.m.
-Designed for partitioned MATPOWER cases, it requires MATLAB R2020a or newer with MATPOWER and YALMIP installed.
-Upon execution, the user provides the partitioned .mat file, ADMM penalty parameter $\rho$, convergence threshold, maximum number of iterations, and optionally the centralized OPF cost for benchmarking via optimality gap.
-After receiving user inputs, the code loads all regional and tie-line data, augments each region’s branch matrix with inter-regional tie-lines, and initializes bus voltage angle variables and indexing.
-The core ADMM loop then solves each region’s DC OPF using YALMIP, minimizing generation cost under local constraints (e.g., power balance, generator limits) while adding consistency penalties on shared tie-line angles as modeled in (1).
-A helper function handles the regional problem formulation and solution.
-After each iteration, the code evaluates tie-line angle consistency and updates dual variables accordingly.
-Primal residuals are recorded, and convergence is declared if the maximum error drops below the user-defined threshold; otherwise, the process continues until the iteration limit.
-The optimality gap is computed and plotted alongside residuals if a centralized cost is provided.
-Progress and convergence status are printed in the command window.
-The DPLib DC OPF solver supports any number of regions and is tailored for use with partitioned .mat files, enabling reproducible distributed OPF testing. Algorithm 2 summarizes its workflow.
-Algorithm 2 DPLib OPF solver
-1:Input: Partitioned MATPOWER data, ADMM parameters $\rho$, convergence tolerance $\epsilon$, maximum iterations
-2: Load the partitioned case data for all regions
-3: Augment each region with its corresponding tie-line data
-4: Initialize region variables including local decision variables and dual variables
-5: Retrieve ADMM parameters and consistency constraints
-6: Identify and prepare interface buses at region boundaries
-7:for each ADMM iteration do
-8:  Solve local OPF for each region independently
-9:  Update dual variables based on consistency violations at bordering buses
-10:  if convergence criteria are met then
-11:   break
-12:  end if
-13:end for
-14: Visualize and save final solution
+The DPLib DC OPF module implements a distributed DC OPF solver based on the ADMM.
+The MATLAB implementation, available in the file DistDCOPF.m and publicly accessible on GitHub, is designed for partitioned MATPOWER cases and supports the YALMIP optimization backends in MATLAB.
+The solver requires MATLAB R2020a or newer with MATPOWER installed; YALMIP must also be added to the MATLAB path.
+At execution time, the user provides the partitioned MATPOWER file, the ADMM penalty parameter rho, a convergence tolerance, a maximum number of iterations, a $\rho_{\mathrm{DC}}^{\max}$, and optionally the centralized OPF cost for computing the optimality gap.
+YALMIP emphasizes modeling clarity and solver flexibility, and is suitable for rapid prototyping and debugging.
+After receiving the user inputs, DistDCOPF.m loads all regional data, reconstructs inter-regional tie-lines, and augments each region’s branch matrix accordingly.
+Bus-angle variables, tie-line indices, and interface bus mappings are then initialized, along with the ADMM dual variables.
+During each ADMM iteration, every region solves its local DC OPF problem independently using the selected backend.
+The local objective minimizes the generation cost subject to regional constraints and tie-line consistency penalties.
+Following all local solves, the solver computes the angle mismatches on shared tie-lines, updates the dual variables, records the maximum primal residual, and checks the convergence condition.
+If a centralized optimal cost is available, the optimality gap is also computed and plotted.
+The procedure continues until convergence is reached or the iteration limit is met.
+The solver supports any number of regions and enables reproducible distributed OPF studies. Its overall workflow is summarized in Algorithm 2.
+Algorithm 2 DPLib Distributed OPF Solver
+1: Input partitioned MATPOWER data, ADMM parameters rho, tolerance epsilon, maximum iterations
+2: Load all regional data and global tie-line table
+3: Augment each region with its corresponding inter-regional tie-line rows
+4: Initialize regional variables, interface indices, and ADMM dual variables
+5: Prepare regional consistency constraints and interface bus mappings
+6:for each ADMM iteration do
+7:  for each region do
+8:   Solve regional OPF
+9:  end for
+10:  Compute tie-line angle mismatches across all region boundaries
+11:  Update dual variables based on consistency violations
+12:  Evaluate primal residual and check convergence criterion
+13:  if convergence achieved then
+14:   Break
+15:  end if
+16:end for
+17: Visualize residuals, optimality gap (if applicable), and final regional solutions
 III-B2 DPLib AC OPF Solver
-The DPLib AC OPF solver implements a distributed ADMM-based solver in MATLAB for any power system preprocessed using the DPLib partitioning tool.
-It enables large-scale distributed analysis, where each region independently solves its local AC OPF while coordinating inter-regional tie-lines.
-The solver developed in this work is publicly available at GitHub, including the MATLAB implementation file named DistACOPF.m.
-To run the code, the user must have MATLAB R2020a or later, with MATPOWER and IPOPT properly configured.
-Upon launch, the code prompts for the partitioned .mat file, ADMM penalty parameter $\rho$, convergence threshold, maximum number of iterations, and optionally the centralized OPF cost for benchmarking via optimality gap.
-After loading the partitioned data, the code augments each region’s branch matrix with its tie-lines and initializes region-specific variables (e.g., bus counts, initial voltage values).
-The main ADMM loop iteratively solves each region’s local AC OPF via the helper function Acopf, based on (2), updating voltage magnitudes and angles.
-Dual variables are adjusted using consistency gradients to enforce agreement across tie-lines.
-Convergence is declared if the worst primal residual falls below the threshold.
-The optimality gap is computed per iteration if a centralized cost is provided.
-Final plots show convergence trends, and status messages are printed throughout.
-The overall process flow of the DPLib AC OPF solver is presented in algorithm 2.
+The DPLib AC OPF solver provides a MATLAB implementation of a distributed ADMM-based AC OPF algorithm that operates on any power system previously partitioned using the DPLib framework.
+The solver is implemented in the publicly available file DistACOPF.m, accessible through the GitHub repository at GitHub.
+The code requires MATLAB R2020a or later, along with a properly configured installation of MATPOWER and IPOPT.
+The function begins by requesting the partitioned .mat file and the user-defined ADMM parameters, including the penalty value $\rho$, the convergence tolerance, the maximum number of iterations, the $\rho_{\mathrm{AC}}^{\max}$, and optionally the centralized AC OPF cost for computing the optimality gap.
+After these inputs are provided, the solver loads the partitioned data and augments the branch matrices of each region with their associated tie-lines.
+It then initializes regional variables describing bus counts, voltage magnitudes and angles, and other quantities needed to enable the local AC OPF solves and the coordination steps.
+The main ADMM loop is implemented within the DistACOPF.m function.
+In each iteration, every region independently solves its local nonlinear AC OPF using the helper function acopf, which is based on the formulation in equation (4).
+The updated voltage magnitudes and phase angles from each region are then exchanged virtually through the consistency constraints applied on the tie-lines.
+The dual variables are modified at each step using the mismatch in the angle and magnitude variables across the interfaces, ensuring that neighboring regions gradually reach agreement.
+The loop continues until the maximum primal residual across all tie-lines drops below the specified tolerance.
+If a centralized AC OPF cost is provided, the function also evaluates the optimality gap at every iteration to quantify the progress toward the global optimum.
+Once the iterations terminate, the solver prints summary information and generates plots that illustrate the convergence behavior of the primal residual and, when applicable, the optimality gap.
+The complete algorithmic flow implemented in DistACOPF.m is summarized in Algorithm 2.
 IV DPLib: Benchmark Test Cases
-Distributed versions of many widely used MATPOWER benchmark systems are generated.
-Each test case is processed using the DPLib graph-based partitioning tool and tested using the DPLib ADMM-based DC and AC OPF solvers described earlier.
-Across all systems, both solvers successfully converge to feasible solutions.
-All partitioned cases and solver outputs are publicly available as part of the DPLib at GitHub.
-In the following subsections, we explain some of the distributed test cases generated and tested as benchmarks in DPLib.
-In addition to evaluating convergence using the worst-case primal residual and the optimality gap, further validation tests are performed on the generated multi-region systems to verify their consistency with physical laws and operational constraints of the power grid.
-Specifically, we verified that upon the convergence of the ADMM iterations, the total power generation across all regions closely matches the total system demand, accounting for minor discrepancies due to modeling approximations and convergence thresholds.
-Furthermore, we confirmed that active power flows across each inter-regional tie-line are consistent from both sides of the interface.
-All tested datasets successfully passed these validation checks, further demonstrating the correctness and reliability of the generated partitioned cases.
-To facilitate evaluation of the benchmark test cases using OPF solvers, all primal variables were initialized with cold-start conditions: voltage angles were set to zero and voltage magnitudes to one per unit.
-The dual variables, corresponding to Lagrange multipliers associated with consistency constraints, were initialized to one.
-Penalty parameters $\rho$ in the ADMM-based solvers were selected case by case to ensure convergence.
-IV-A Case: pglib_opf_case200_tamu
-The pglib_opf_case200_activ system is a synthetic 200-bus benchmark from the IEEE PES Power Grid Library for optimal power flow studies.
-Geographically situated in the central region of Illinois, this case models a 230/115 kV transmission network designed to serve a load distribution that mimics the population profile of its geographic footprint.
-The transmission system was generated following the methods outlined in [19], ensuring statistical similarity to real-world grids without modeling any actual transmission lines.
-Developed as part of the U.S.
-ARPA-E GRID DATA research project, the system contains no critical energy infrastructure information and was created using PowerWorld Simulator, then converted to MATPOWER format using version 6.
-The dataset is publicly available through PGLib and licensed under the Creative Commons Attribution 4.0 International License.
-The pglib_opf_case200_tamu test case was partitioned using our graph-based spectral clustering method, which ensured full bus inclusion and topological connectivity.
-The best clustering configuration identified by the algorithm consisted of three electrically coherent regions, interconnected through 10 tie-lines, within 10 iterations.
-During the partitioning, the reference bus, bus 189, was assigned to Region 2 (mpc_regionR2).
-No buses were omitted, and the total number of buses across all regions matches the original MATPOWER dataset, with 200 buses, 245 branches, and 49 generators.
-The consistency of branch count, generator count, and overall system topology confirms the fidelity of the partitioning process.
-The final multi-region topology has been visualized and is shown in Fig.
-2.
-The corresponding partitioned files are saved as pglib_opf_case200_tamu_3regions.mat and pglib_opf_case200_tamu_3regions.m.
-Both distributed DC OPF and AC OPF solvers have been executed on this partitioned data, and their convergence trends are illustrated through the worst primal residuals presented in Fig.
-3 with $\rho=1000$.
-Figure 2: Topology of pglib_opf_case200_tamu partitioned into three regions with 10 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 3: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case200_tamu.
-IV-B Case: pglib_opf_case300_ieee
-The pglib_opf_case300_ieee system is a widely used benchmark from the IEEE PES Power Grid Library for optimal power flow studies.
-This test case represents the IEEE 300-bus system, which was initially provided in the IEEE Common Data Format and converted into MATPOWER format on September 20, 2004, using the cdf2matp utility.
-The dataset was sourced from the University of Washington’s power systems test case archive.
-The system corresponds to typical operational conditions, with a base power of 100 MVA and includes detailed modeling of generation, loads, and transmission elements.
-Using our graph-based spectral clustering approach, we have partitioned the pglib_opf_case300_ieee system into three electrically coherent regions.
-Before clustering, the original bus numbering in the MATPOWER file was reordered to ensure consistent indexing.
-The graph topology of the network was verified to be fully connected, allowing for a valid partitioning process.
-The algorithm identified the best clustering configuration in 10 iterations, resulting in 3 regions interconnected through 12 tie-lines.
-The system’s slack bus, originally defined at bus 257, was successfully assigned to Region 1 (denoted as mpc_regionR1).
-After partitioning, the integrity of the dataset was confirmed: all 300 buses, 411 branches, and 69 generators were fully included in the regional breakdown, with no data omitted.
-The final partitioned system has been saved as both a .mat file and a script-based .m file named pglib_opf_case300_ieee_3regions.
-A visual representation of the multi-regional topology is shown in Fig.
-4.
-The distributed DC OPF and AC OPF solvers were successfully executed on this dataset, and the convergence behavior is demonstrated through the worst primal residual plots shown in Figs.
-5(a) and 5(b) with $\rho=1000$.
-Figure 4: Topology of pglib_opf_case300_ieee partitioned into three regions with 12 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 5: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case300_ieee.
-IV-C Case: pglib_opf_case500_tamu
-The pglib_opf_case500_tamu case is a synthetic 500-bus power system in the IEEE PES Power Grid Library for OPF studies.
-This case was developed as part of the ARPA-E GRID DATA research initiative and is geographically based in the northwestern region of South Carolina.
-It incorporates a transmission network operating at 345 kV and 138 kV voltage levels, constructed to supply power to a synthetic load distribution that mirrors the population density of the modeled region.
-The system does not represent any real grid infrastructure but is created using statistical modeling techniques described in [19], ensuring structural similarities to real transmission systems while maintaining confidentiality.
-Using our graph-based spectral clustering framework, the pglib_opf_case500_tamu benchmark system was partitioned into five electrically coherent regions.
-Before clustering, the bus numbers from the original MATPOWER file were analyzed to preserve consistent indexing.
-The graph representation of the system was confirmed to be fully connected, ensuring the validity of the partitioning procedure.
-The spectral clustering algorithm converged in 10 iterations and identified an optimal structure of 5 regions interconnected through 16 tie-lines.
-The reference bus, originally defined as bus 17, was successfully assigned to Region 2 (denoted as mpc_regionR2).
-Post-partitioning checks confirmed that all 500 buses, 597 branches, and 90 generators were preserved without omission or duplication.
-The partitioned data has been saved in both .mat and script-based .m formats under the name pglib_opf_case500_tamu_5regions.
-A visual illustration of the multi-regional layout is presented in Fig.
+TABLE I: Partitioning summary with centralized DC/AC OPF costs.
+System Size Partition Penalty Param. Centralized OPF Cost
+Case Buses Branches Gens Regions Tie-lines Slack $\rho$ $\rho_{\mathrm{DC}}^{\max}$ $\rho_{\mathrm{AC}}^{\max}$ DC AC
+pglib_opf_case200_tamu $200$ $245$ $49$ 33 $10$ R2 11 1e81108 1e81108 $34,653$ $27,558$
+pglib_opf_case300_ieee $300$ $411$ $69$ 33 $12$ R1 11 1e81108 1e81108 $517,310$ $565,220$
+pglib_opf_case500_tamu $500$ $597$ $90$ 55 $16$ R2 11 1e81108 1e81108 $91,841$ $72,578$
+pglib_opf_case1354_pegase $1,354$ $1,991$ $260$ 88 $62$ R1 11 1e81108 1e81108 $1,218,028$ $1,258,844$
+pglib_opf_case2869_pegase $2,869$ $4,582$ $510$ $12$ $95$ R6 11 1e51105 1e81108 $2,386,170$ $2,462,790$
+pglib_opf_case4661_sdet $4,661$ $5,997$ $1,176$ $14$ $284$ R3 11 1e61106 1e61106 $2,217,130$ $2,255,540$
+pglib_opf_case9241_pegase $9,241$ $16,049$ $1,445$ $18$ $197$ R17 11 1e61106 1e51105 $6,044,174$ $6,243,090$
+This section presents a detailed description and verification of several DPLib benchmark systems and their corresponding multi-region topologies.
+All test cases originate from the PGLib-OPF library and are partitioned using the spectral clustering procedure of Algorithm 1 based on the unweighted Laplacian matrix.
+Before partitioning, every system undergoes bus renumbering when required, connectivity verification, slack-region assignment, and integrity checks to ensure that all buses, branches, and generators are preserved.
+The resulting number of regions, tie-lines, and solver parameters are summarized in Table I.
+All region-labeled MATPOWER files, verification codes, and the topology and verification figures included here are available in the DPLib GitHub repository.
+To make the library easy to use and to avoid case-by-case tuning, the ADMM penalty parameters are initialized to $\rho^{0}=1$ for all benchmarks in both the distributed DC-OPF and AC-OPF solvers, after which $\rho$ is adjusted automatically via the adaptive residual-balancing rule where $\tau_{\mathrm{incr}}=2,\tau_{\mathrm{decr}}=0.5$.
+In the distributed DC-OPF solver, the adaptive update parameters are set internally to
+$\mu=\max\!\left(1,\ \frac{15}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right),$ (8)
+where $n_{\mathrm{reg}}$ denotes the number of regions. For the distributed AC-OPF solver, the same heuristic is applied separately to the angle and voltage-magnitude channels using
+$\mu_{\theta}=\max\!\left(1,\ \frac{15}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right),\;\mu_{v}=\max\!\left(1,\ \frac{10}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right).$ (9)
+These choices scale the residual-balance thresholds with the partition granularity while keeping fixed and interpretable increase/decrease factors, allowing the adaptive-$\rho$ mechanism to operate robustly across different network sizes and region counts with minimal user intervention.
+The verification procedure is implemented in a MATLAB script that processes partitioned cases in a unified workflow.
+For each system, the script first loads the corresponding MATPOWER case and calls the centralized solvers run_dcopf_centralized and run_acopf_centralized.
+These functions compute reference DC OPF and AC OPF objective values and return solver status information.
+The centralized costs serve as benchmarking values to compute the optimality gap of the distributed solutions if the centralized DC OPF is feasible.
+The distributed DC OPF solver is run using the function DistDCOPF with prescribed tolerance and maximum iteration count.
+Similarly, the distributed AC OPF solver DistACOPF is called following the structure of Algorithm 2.
+For each distributed run, the script tracks the worst primal residual across tie-lines and the optimality gap with respect to the centralized objective value.
+Table I reports placeholder centralized DC and AC objective values for all systems.
+In the numerical experiments, for all cases where the centralized solvers return feasible solutions, the distributed solvers drive the primal residual down to the threshold and achieve an optimality gap below one percent, indicating that the distributed solutions are both feasible and near-optimal.
+Case pglib_opf_case200_tamu
+The system pglib_opf_case200_tamu is a 200-bus synthetic model representing a 230/115 kV network in Illinois.
+It was generated as part of the ARPA-E GRID DATA project and is designed to capture realistic transmission characteristics and operational patterns.
+The system is partitioned into three regions connected through ten tie-lines.
+The resulting multi-region topology is shown in Fig.
+3.
+Figure 3: Three-region topology for case200_tamu.
+Figures 4 and 5 report the DC and AC primal residuals and optimality gaps for pglib_opf_case200_tamu.
+Figure 4: primal residual (left) and optimality gap (right)
+Figure 5: primal residual (left) and optimality gap (right)
+Case pglib_opf_case300_ieee
+The system pglib_opf_case300_ieee corresponds to the well-known IEEE 300-bus test case.
+Originally formulated in the IEEE Common Data Format and subsequently converted to MATPOWER, it has long served as a standard benchmark for power system optimization studies.
+The system is divided into three regions, with interconnections distributed across twelve tie-lines.
+The extracted topology is shown in Fig.
 6.
-Distributed DC OPF and AC OPF solvers were successfully executed for this case, and the convergence profiles are illustrated by the worst primal residual curves in Figs.
-7(a) and 7(b) with $\rho=1000$.
-Figure 6: Topology of pglib_opf_case500_tamu partitioned into five regions with 16 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 7: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case500_tamu.
-IV-D Case: pglib_opf_case1354_pegase
-The pglib_opf_case1354_pegase dataset is a detailed test case that emulates the size and complexity of the European high-voltage transmission network.
-It includes 1,354 buses, 260 generators, and 1,991 branches operating at 220 and 380 kV voltage levels.
-While the data is fictitious and does not reflect real-world infrastructure, it is highly suitable for validating power system optimization tools and techniques.
-This benchmark was developed as part of the Pan European Grid Advanced Simulation and State Estimation (PEGASE) project.
-It incorporates unique modeling features such as asymmetric shunt elements and generator aggregations.
-The pglib_opf_case1354_pegase benchmark system was partitioned into eight electrically coherent regions using our graph-based spectral clustering method.
-The original bus indices from the MATPOWER file were reordered to ensure consistency during partitioning.
-The graph representation of the power network was verified to be fully connected, ensuring that a valid and complete partitioning could be achieved.
-The clustering algorithm successfully identified an optimal partitioning configuration within 10 iterations, resulting in 8 regions interconnected through 62 tie-lines.
-The system’s reference bus, which is located at bus 640 in the original data, was assigned to Region 1 (denoted as mpc_regionR1).
-The integrity of the data was verified after partitioning: all 1354 buses, 1991 branches, and 260 generators were accurately distributed across the regions with no missing components.
-The partitioned data files were saved in both .mat and .m formats under the name pglib_opf_case1354_pegase_8regions.
-A graphical representation of the resulting multi-regional topology is provided in Fig.
-8.
-Both distributed DC OPF and AC OPF solvers were applied to this case study, with convergence behavior illustrated via worst primal residual plots in Figs.
-9(a) and 9(b) with $\rho=1000$.
-Figure 8: Topology of pglib_opf_case1354_pegase partitioned into eight regions with 62 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 9: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case1354_pegase.
-IV-E Case: pglib_opf_case2383wp_k
-The pglib_opf_case2383wp_k dataset models the Polish national transmission grid under peak load conditions during the winter of 1999–2000.
-It includes a comprehensive representation of 400, 220, and 110 kV networks and originates from a reduced version of the broader 7500+ bus European UCTE system.
-To simplify the model while preserving essential characteristics, tie lines connecting to neighboring countries were replaced by artificial load or generator buses.
-Generator aggregation and cost approximations were applied to reflect dispatchability in the Polish energy market.
-Using the proposed graph-based spectral clustering method, the pglib_opf_case2383wp_k benchmark system was partitioned into twelve electrically meaningful regions.
-The underlying graph structure of the power system was verified to be fully connected, confirming the feasibility of multi-regional partitioning.
-The algorithm determined the optimal clustering within 10 iterations, producing 12 regions linked by 99 inter-regional tie-lines.
-The system’s slack bus, originally defined at bus 18 in the MATPOWER case, was assigned to Region 4 (referred to as mpc_regionR4).
-Following partitioning, a full consistency check verified that all 2383 buses, 2896 branches, and 327 generators were preserved and correctly distributed across the regional datasets.
-No components were omitted during the process.
-The results were stored as both a .mat file and an executable .m script, saved under the name pglib_opf_case2383wp_k_12regions.
-The partitioned network topology is visually represented in Fig.
-10.
-The distributed DC OPF and AC OPF solvers are implemented and their convergence behavior is shown through the worst primal residual plots in Figs.
-11(a) and 11(b) with $\rho=2000$.
-Figure 10: Topology of pglib_opf_case2383wp_k partitioned into twelve regions with 99 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 11: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case2383wp_k.
-IV-F Case: pglib_opf_case3012wp_k
-The pglib_opf_case3012wp_k dataset provides a detailed model of the Polish transmission grid during the winter 2007–2008 evening peak demand period.
-It represents the operational state of the 400, 220, and 110 kV networks and is a high-fidelity snapshot of the Polish system during that time.
-As part of the Power Grid Library benchmark suite, it is widely used for testing and validating optimization algorithms in realistic, large-scale scenarios.
-Using the proposed graph-based spectral clustering framework, the pglib_opf_case3012wp_k benchmark system was partitioned into twelve electrically coherent regions.
-Before partitioning, the original MATPOWER bus numbers were reordered to enable consistent indexing throughout the graph representation.
-The full graph connectivity was verified to ensure that all buses could be properly included in the clustering process.
-Within 10 iterations, the partitioning algorithm identified an optimal configuration that divides the system into 12 regions connected by 100 inter-regional tie-lines.
-The slack bus, located initially at bus 37 in the MATPOWER data, was assigned to Region 12 (denoted as mpc_regionR12).
-Following the partitioning process, a verification step confirmed that the distributed representation included all 3012 buses, 3572 branches, and 502 generators.
-These components were preserved without omission or duplication.
-The resulting data was saved both as a .mat file and as a script-based .m file under the name pglib_opf_case3012wp_k_12regions.
-A visual overview of the final partitioned topology is shown in Fig.
+Figure 6: Three-region topology for case300_ieee.
+Figures 7 and 8 show the DC and AC verification results for pglib_opf_case300_ieee.
+Figure 7: primal residual (left) and optimality gap (right)
+Figure 8: primal residual (left) and optimality gap (right)
+Case pglib_opf_case500_tamu
+The system pglib_opf_case500_tamu is a 500-bus synthetic model of the South Carolina transmission grid, also produced under the ARPA-E GRID DATA initiative.
+It contains detailed line, transformer, and generator data reflecting real-world system behavior.
+The network is partitioned into five regions connected by sixteen tie-lines.
+The corresponding topology appears in Fig.
+9.
+Figure 9: Five-region topology for case500_tamu.
+For pglib_opf_case500_tamu, Figures 10 and 11 illustrate DC and AC primal residuals and optimality gaps, respectively.
+Figure 10: primal residual (left) and optimality gap (right)
+Figure 11: primal residual (left) and optimality gap (right)
+Case pglib_opf_case1354_pegase
+The dataset pglib_opf_case1354_pegase is part of the PEGASE project, which produced large-scale synthetic European transmission networks intended for stability and optimization studies.
+This 1354-bus system includes high-density meshed areas characteristic of continental European grids.
+It is partitioned into eight regions with numerous inter-region connections reflecting the underlying network complexity.
+The multi-region layout is shown in Fig.
 12.
-Using this distributed dataset, the DC OPF and AC OPF solvers were executed successfully, with their convergence behavior illustrated by the worst primal residual plots shown in Figs.
-13(a) and 13(b) with $\rho=2000$.
-Figure 12: Topology of pglib_opf_case3012wp_k partitioned into twelve regions with 100 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 13: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case3012wp_k.
-IV-G Case: pglib_opf_case4661_sdet
-The pglib_opf_case4661_sdet dataset is a large-scale synthetic power system model with 4,661 buses.
-It was created using the Sustainable Data Evolution Technology tool and is publicly available through the eGRID Data portal.
-This model is part of the PGLib OPF v19.05 benchmark suite and is provided in the public domain to support research in scalable and realistic power system analysis.
-The pglib_opf_case4661_sdet benchmark system was partitioned into fourteen electrically coherent regions using the proposed graph-based spectral clustering methodology.
-Before partitioning, the original bus numbering in the MATPOWER data was reordered to enable consistent and systematic region assignment.
-The underlying graph representation of the network was verified to be fully connected, ensuring that all components were eligible for inclusion in the partitioning process.
-After 10 iterations, the algorithm identified an optimal clustering configuration with 14 regions and 284 inter-regional tie-lines.
-The reference (slack) bus, originally located at bus 2625 in the MATPOWER data, was assigned to Region 3 (denoted as mpc_regionR3).
-A validation step was performed after partitioning to confirm the structural integrity of the distributed dataset.
-It verified that all 4661 buses, 5997 branches, and 1176 generators from the original MATPOWER case were successfully preserved and assigned across the generated regional files.
-The partitioned data has been saved in both .mat and script-based .m formats under the name pglib_opf_case4661_sdet_14regions.
-A visual summary of the multi-regional topology is presented in Fig.
-14.
-Subsequently, the distributed DC OPF and AC OPF solvers were executed on the partitioned system, and their convergence characteristics are shown through the worst primal residual plots in Figs.
-15(a) and 15(b) with $\rho=4000$.
-Figure 14: Topology of pglib_opf_case4661_sdet partitioned into fourteen regions with 284 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 15: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case4661_sdet.
-IV-H Case: pglib_opf_case6468_rte
-The pglib_opf_case6468_rte [20] dataset provides a realistic snapshot of the French high voltage and very high voltage transmission network in 2013.
-It contains 6,468 buses, 1,295 generators, and 9,000 branches, offering a highly detailed and complex test case for power system optimization.
-While the data are based on real grid topologies, they are intended solely for academic and research purposes and should not be used for operational or planning decisions related to the French or European power grids.
-This dataset is part of the IEEE PES Power Grid Library (PGLib-OPF v19.05) and was developed by Réseau de Transport d’Électricité (RTE), the French transmission system operator.
-It supports validating optimization algorithms, mathematical modeling techniques, and simulation tools in large-scale power system analysis.
-The pglib_opf_case6468_rte benchmark was successfully partitioned into electrically coherent regions using our spectral clustering-based partitioning algorithm.
-This method ensured that the network’s graph structure remained fully connected.
-The best clustering outcome was identified within 10 iterations, resulting in 16 regions interconnected by 204 inter-regional tie-lines.
-Bus renumbering was applied to the original MATPOWER case to enable consistent indexing across the partitioning workflow.
-The system’s slack bus, labeled as bus 4733, was assigned to Region 6 (mpc_regionR6).
-All buses, branches, and generators from the original MATPOWER dataset were accounted for in the partitioned output, confirming the completeness and validity of the partitioning.
-Specifically, the resulting distributed dataset includes all 6468 buses, 9000 branches, and 1295 generators from the original case.
-The partitioned version has been saved in both .mat and script-based .m formats under the name pglib_opf_case6468_rte_16regions.
-A visualization of the region-level partitioned topology is presented in Fig.
-16.
-To validate the partition, we applied the distributed DC OPF and AC OPF solvers, which successfully converged.
-Their worst primal residual plots over ADMM iterations are shown in Fig.
-17 with $\rho=6000$.
-Figure 16: Topology of pglib_opf_case6468_rte partitioned into 16 regions with 204 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 17: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case6468_rte.
-IV-I Case: pglib_opf_case9241_pegase
-The pglib_opf_case9241_pegase dataset provides a highly detailed representation of the European high voltage transmission network.
-It includes 9,241 buses, 1,445 generators, and 16,049 branches operating across multiple voltage levels, including 750, 400, 380, 330, 220, 154, 150, 120, and 110 kV.
-Although the dataset is fictitious and does not reflect real-world data, it is valuable for validating optimization algorithms and simulation tools in large-scale power system research.
-The data were released as part of the Pan European Grid Advanced Simulation and State Estimation (PEGASE) project [21].
-The pglib_opf_case9241_pegase benchmark represents one of the largest systems in the MATPOWER repository and was successfully partitioned using our graph-based spectral clustering methodology.
-The partitioning process ensured complete inclusion of all 9241 buses and confirmed the full connectivity of the underlying power grid graph.
-The best clustering configuration was determined within 10 iterations, resulting in 18 regions connected by 197 inter-regional tie-lines.
-The system’s reference bus, numbered 4231, was assigned to Region 17 (mpc_regionR17).
-The reorganization of the data preserves the integrity of the original MATPOWER structure, with a total of 16049 branches and 1445 generators fully accounted for across the distributed regions.
-The partitioned results were saved in both .mat and .m formats under the name pglib_opf_case9241_pegase_18regions.
-A visual representation of the multi-region topology is provided in Fig.
+Figure 12: Eight-region topology for case1354_pegase.
+Figures 13 and 14 present the DC and AC verification curves for pglib_opf_case1354_pegase.
+Figure 13: primal residual (left) and optimality gap (right)
+Figure 14: primal residual (left) and optimality gap (right)
+Case pglib_opf_case2869_pegase
+The system pglib_opf_case2869_pegase represents a large portion of the European transmission network, consisting of 2869 buses, 510 generators, and 4582 branches.
+The partitioning procedure divides the system into twelve regions, interconnected through ninety-five tie-lines.
+The resulting multi-region topology is shown in Fig.
+15.
+All buses, generators, and branches are fully accounted for across the regions, and the reference bus (Bus 1314) is assigned to Region 6.
+Figure 15: Twelve-region topology for case2869_pegase.
+For pglib_opf_case2869_pegase, Figures 16 and 17 illustrate that both DC and AC primal residuals converge to the threshold, while the optimality gaps remain below one percent.
+Figure 16: Primal residual (left) and optimality gap (right)
+Figure 17: Primal residual (left) and optimality gap (right)
+Case pglib_opf_case4661_sdet
+The system pglib_opf_case4661_sdet is a large-scale synthetic model developed using the Sustainable Data Evolution Technology (SDET) framework.
+With 4661 buses and dense connectivity, it reflects realistic large-area transmission dynamics.
+The system is separated into fourteen regions connected by 284 tie-lines.
+The partitioned structure is shown in Fig.
 18.
-This dataset has been validated by executing the distributed DC OPF and AC OPF solvers, with the convergence of the worst primal residuals depicted in Fig.
-19 with $\rho=8000$.
-Figure 18: Topology of pglib_opf_case9241_pegase partitioned into 18 regions with 197 tie-lines.
-(a) DC OPF
-(b) AC OPF
-Figure 19: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case9241_pegase.
+Figure 18: Fourteen-region topology for case4661_sdet.
+For pglib_opf_case4661_sdet, Figures 19 and 20 indicate that the DC and AC primal residuals converge to the threshold, with optimality gaps remaining below one percent.
+Figure 19: primal residual (left) and optimality gap (right)
+Figure 20: primal residual (left) and optimality gap (right)
+Case pglib_opf_case9241_pegase
+The largest benchmark system considered in DPLib is pglib_opf_case9241_pegase, a synthetic European transmission network derived from the PEGASE project.
+With 9241 buses and more than sixteen thousand branches, it reflects the characteristics of a highly interconnected continental-scale system.
+The partitioning procedure yields eighteen regions linked by 197 tie-lines.
+The resulting multi-region topology is shown in Fig.
+21.
+Figure 21: Eighteen-region topology for case9241_pegase.
+For pglib_opf_case9241_pegase, Figures 22 and 23 confirm that the DC and AC primal residuals converge to the threshold, with optimality gaps remaining within one percent, even for this largest benchmark.
+Figure 22: primal residual (left) and optimality gap (right)
+Figure 23: primal residual (left) and optimality gap (right)
 V Conclusion
@@
 Numerical results demonstrate the accuracy of the generated test cases and convergence of the solvers, making DPLib a practical foundation for future research in scalable, privacy-preserving, and resilient power system analysis.
+To support long-term usability, we are committed to maintaining DPLib as an actively developed open-source project.
+Future updates will include bug fixes, feature enhancements, expanded test systems, and improvements to the distributed OPF solvers as new algorithms become available.
+The DPLib repository will continue to be updated and versioned through GitHub, where users may report issues, request features, and contribute to the codebase.